### v1

# Euler-type approximation for the invariant measure: An abstract framework

Abstract.
We establish a general framework to study the rate of convergence of a Euler type approximation scheme with decreasing time steps to the invariant measure, for a general class of stochastic systems. The error is measured in general Wasserstein distances, which enables to encompass cases with non global contractivity conditions. Our main assumption is a coupling property which is expressed in terms of the one-step approximation. We show that the proposed set-up can be applied to a wide range of equations that may be law dependent, such as Langevin equations, reflected equations, Boltzmann type equations and for a recent McKean Vlasov type model for neuronal activity.
Key words and phrases:
Invariant measure approximation, Euler-type scheme with decreasing step, McKean-Vlasov equations, Boltzmann equations
1991 Mathematics Subject Classification:
37M25, 60G99, G5C99
1. Introduction
The problem of existence and uniqueness of the invariant measure for a Markov semigroup has a long history. We refer the reader to the influential paper by Meyn and Tweedie [MT93] and Hairer’s lecture notes [Hai21]. In direct connection to this question, the speed of convergence of the semigroup to the invariant measure has been also widely studied, see in particular Bakry et al. [BCG08], Cattiaux Guillin [CG09], Hairer [Hai21] and the references therein. Unless the semigroup can be simulated exactly, one typically uses approximation schemes of the semigroup to then approximate the invariant measure. Approximation schemes for the invariant measure associated to some stochastic process are thus built in two steps: a first step deals with the error between the invariant measure and the distribution of the semigroup for large times, and second step deals with the error between the semigroup and its approximation.
In this sense, approximations based on Euler type schemes has also been an important research topic in the past years, particularly in the specific framework of the Langevin diffusion that is related to gradient descent algorithms that are widely used in Statistics and Machine Learning. We may distinguish two streams of research. The first one considers a constant time step for the Euler type approximation and then analyses the error between the invariant measures of the continuous and discretised semigroup, see in particular Talay [Tal02], Mattingly et al. [MSH02], Dalalyan and Tsybakov [DT12], Dalalyan [Dal17], Durmus and Moulines [DM17], Brosse et al. [BDMS19], Crisan et al. [CDO21] and Chen et al. [CDSX23]. The second stream of research, initiated by Lamberton and Pagès [LP02, LP03] and Lemaire [Lem05] considers instead an approximation scheme with a decreasing sequence of time steps, which allows to get directly the convergence to the invariant measure of the continuous semigroup. This point of view has been used in several recent works: see Pagès and Panloup [PP23], Bally and Qin [BQ24].
Furthermore a number of stochastic equations in which the law of the solution is part of the equation, such as in the cases of McKean-Vlasov or Boltzmann type equations, do not seem to be within the framework of Markov semigroups. The literature about the approximation of the invariant measure for such systems is less developed but very active in the last years (see Schuh ans Souttar [SS24], Cormier [Cor24] and Du et. al. [DJL23] and the references therein). Having this motivation in mind, we focus on the approximations to the invariant measures for flows of transformations on the Wasserstein space, indexed on time. First, under appropriate hypotheses, we obtain exponential speed of convergence to the invariant measure as the time goes to infinity. Second, considering an approximation scheme with a decreasing step, we obtain in a general framework the rate of convergence in Wasserstein distance between the invariant measure and its approximation. Our work is related to the recent paper of Schuh and Souttar [SS24] who provide also a general framework to get uniform in time estimates between the continuous time process and its approximation, but for a constant time step approximation. As a consequence, there are significant differences between their framework and ours that we present in detail in Remark 2.13 below.
Let us be more precise and present the main results of the paper. In the first section (Section 2) we introduce the main objects. We consider a separable Polish metric space $(B,d)$. We denote $\mathcal{P}(B)$ the space of probability measures on $B$ and, for $p\geq 1$, $\mathcal{P}_{p}(B)$ is the set of probability measures with finite $p$ moment, endowed with the Wasserstein distance $W_{p}$. The main object of study in our paper is a family of transformations $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$. Our first problem, is to give conditions that ensure the existence and uniqueness of the invariant probability measure, that is a probability measure $\nu$ such that $\Theta_{s,t}(\nu)=\nu$ for all $s<t$.
The next step towards our main goal is to define the approximation scheme. With some abuse of notation, we define for a time grid $\pi=\{s=t_{0}<t_{1}<...<t_{n}<$...\} such that $t_{n}\to\infty$,
$\Theta_{t_{0},t_{k}}^{\pi}(\mu)=\Theta_{t_{k-1},t_{k}}\circ\Theta_{t_{k-1},t_{k-2}}\circ...\circ\Theta_{t_{0},t_{1}}.$
In the present case, we have in mind that $\Theta_{t_{i},t_{i+1}}$ represents an approximation, such as the one step Euler scheme between times $t_{i}$ and $t_{i+1}$, and therefore $\Theta_{t_{0},t_{k}}^{\pi}$ represents the corresponding Euler scheme on the time grid $\pi$. Our goal is to give conditions so that $\Theta_{t_{0},t_{k}}^{\pi}$ converges in $p$-Wasserstein distance to $\nu$ as $k\to\infty$ and to determine its rate of convergence.
A first hypothesis which is in force in our approach is a uniform bound on the $p$-moments for some $h>0$:
$M_{p}(\Theta,\mu,h)=\sup_{\pi:t_{i}-t_{i-1}\leq h}\sup_{n\geq 0}\left\|\Theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p}<\infty,$
with the supremum taken on all the partitions $\pi$ and all $n\in{\mathbb{N}}$. In many examples, this property is obtained with a Foster Lyapunov type condition (3.2). In some other examples, this property can be derived from existing results on the corresponding continuous time process.
Next, we introduce a coupling argument that plays a central role in our analysis. We say that two families $\Theta^{1}$ and $\Theta^{2}$ are $(p,b_{*},\varepsilon)$- coupled with $b_{*},\varepsilon>0$ if
$W_{p}^{p}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))\leq(1-(t-s)b_{*})W_{p}^{p}(\mu^{1},\mu^{2})+C(1+\left\|\mu^{1}\right\|_{p}^{p}+\left\|\mu^{2}\right\|_{p}^{p})(t-s)^{1+\varepsilon}$ (1.1)
for 0<ts<h0<t-s<h. Lemma 2.6 proves then under this condition that
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq e^{-b_{*}(t_{n}-t_{0})}W_{p}^{p}(\mu^{1},\mu^{2})+C(1+M_{p}(\Theta^{1},\mu^{1},h)+\left\|\mu^{2}\right\|_{p}^{p})\sigma_{b_{*},\varepsilon}(n)$ (1.2)
with
$\sigma_{b_{*},\varepsilon}(n)=\sum_{k=1}^{n}e^{-b_{*}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}\quad\text{ and }\quad\ \gamma_{k}:=t_{k+1}-t_{k}.$
Following Lamberton and Pagès in [LP02] and [LP03], we get under suitable assumptions on the decreasing time step $\gamma_{n}$ that $\sigma_{b_{*},\varepsilon}(n)\to_{n\to 0}0$ with a precise rate of convergence. A typical example is to take $\gamma_{n}=\frac{1}{h^{-1}+n}$, and we have then $\sigma_{b_{*},\varepsilon}(n)\leq C/n^{\varepsilon}$.
We are now able to present our main abstract results [^1] For exact definitions in the results below, we refer the reader to the main text.. The first one concerns existence and uniqueness of the invariant measure and exponential estimates of the speed of convergence. Let $\theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ be a flow ($\theta_{r,t}\circ\theta_{s,r}=\theta_{s,t}$ for $0\leq s\leq r\leq t$) that is time homogeneous (i.e. $\theta_{s,t}(\mu)=\theta_{0,t-s}(\mu)$) such that $\mu\rightarrow\theta_{s,t}(\mu)$ is continuous for the $W_{p}$ distance. We assume that $\theta$ has a uniform bound on the $p$ moments and moreover that $\theta$ is $(p,b_{*},\varepsilon)$-coupled with itself, meaning that (1.1) holds with $\Theta^{1}=\Theta^{2}=\theta$. Then, there exists a unique invariant probability measure $\nu$ (that is $\theta_{s,t}(\nu)=\nu$ for every $s\leq t$, and one has the following exponential speed of convergence
$\forall\mu\in\mathcal{P}_{p}(B),\quad W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ce^{-(b_{*}\wedge\varepsilon)(t-s)}$ (1.3)
when $b_{*}\not=\varepsilon$ (an extra multiplicative factor (1+(ts))(1+(t-s)) appears when $b_{*}=\varepsilon$). Our second result concerns the approximation of $\nu$ by means of an approximation scheme for the grid $\pi$ with $\gamma_{k}=\frac{1}{h^{-1}+k}$, $k\in\mathbb{N}$. If we suppose in addition to the previous hypotheses that $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ which is $(p,b_{*},\varepsilon)$-coupled with $\theta$ in the sense of (1.1), then for every $\mu\in\mathcal{P}_{p}(B)$
$W_{p}(\Theta_{t_{0},t_{n}}^{\pi}(\mu),\nu)\leq\frac{C}{n^{{\varepsilon\wedge b_{*}}}}.$ (1.4)
Again, this result holds for $b_{*}\not=\varepsilon$ and an extra multiplicative factor $(1+\ln(1+n))$ appears when $b_{*}=\varepsilon$.
In the second part of the paper, we present examples that fit our general framework. The key is to prove the $(p,b_{*},\varepsilon)$ - coupling property (1.2) between the approximation family $\Theta$ and the limit flow $\theta$. In most cases, this is done by using a global contraction property on the coefficients as in Section 5. However, in some specific examples related to the Langevin equation, Eberle [Ebe16] (see also [Sch24]) succeeds in proving the existence of the invariant measure and exponential convergence to equilibrium under a much weaker assumption where the contraction with respect to the Euclidean norm has to be in force outside a compact set only. The argument is based on the construction of a specific distance, which is equivalent to the Euclidean distance, under which strict contraction holds on the whole space. On [Ebe16, Example 1], we prove under the same conditions as the ones stated in [Ebe16] that the Euler scheme converges, and we give as well a rate of convergence. In fact, the formalism developed in our paper allows to use the estimates obtained in [Ebe16] for the flow in order to study the Euler approximation.
Now, we give the main structure of the article. The abstract framework is presented in Section 2 together with our main general results. Section 3 provides interesting tools to fit our general framework. Namely, we show how the sewing lemma can help to get the coupling property (1.2) for $W_{1}$ type distances, and give a Foster-Lyapunov type criterion to get uniform bounds on moments. Then, we give in Section 4 the results concerning the Langevin equation mentioned above. The following Section 5 is devoted to some more examples. In Subsection 5.1, we consider a stochastic equation with reflection driven by a Brownian motion with a $W_{2}$ distance. In Subsection 5.2, we deal with a recent neuronal model which has been introduced in [DMGLP15] and [FL16] and has recently been discussed in [CTV20]. This is a mean field type model driven by a Poisson Point process. We give two results, in terms of $W_{1}$ and $W_{2}$ using different global contraction assumptions in each case. In Subsection 5.3, we consider a Boltzmann type equation - that is an equation driven by a Poisson Point process with intensity depending on the law of the solution. Such equations appear in models of interacting particle systems. We consider two possible regimes: a finite variation regime with the distance $W_{1}$ and a martingale regime with the distance $W_{2}$. The further analysis of the particle systems or of the approximation based on the ergodic measure arising from these law dependent examples will be done in a further research. As many examples consist in doing verification of the required hypotheses of our general framework, they are often similar in nature, and we have put some of them in the additional material in Appendix B, keeping the most illustrative proofs in the main text.
In general, throughout the article, unless constants are written explicitly, they may change from one line to the next even though they are denoted by the same letter $C$.
2. Framework and main results
We introduce first a general framework for convergence that will be applied through out the article.
Let $(B,d)$ be a Polish metric space and $x_{0}\in B$ a given element of $B$. For a probability $\mu$ on $B$ and $p\geq 1$, we denote
$\left\|\mu\right\|_{p}^{p}=\int_{B}d(x,x_{0})^{p}\mu(dx).$ (2.1)
We note $\mathcal{P}(B)$ the space of probability measures on $B$ and $\mathcal{P}_{p}(B)=\{\mu\in{\mathcal{P}}:\|\mu\|_{p}<\infty\}$ the set of probability measures with finite $p$ moment. We endow $\mathcal{P}_{p}(B)$ with the Wasserstein distance $W_{p}$ defined by
$W_{p}^{p}(\mu,\nu)=\inf_{\rho\in\Pi(\mu,\nu)}\int_{B}\int_{B}d(x,y)^{p}\rho(dx,dy),$
where $\Pi(\mu,\nu)=\{\rho\in{\mathcal{P}}(B\times B):\forall A\in\mathcal{B}(B),\ \rho(A\times B)=\mu(A),\rho(B\times A)=\nu(A)\}$ is the set of all probability measures on $B\times B$ with marginals $\mu$ and $\nu$. We recall that $\mathcal{P}_{p}(B)$ is complete with respect to the distance $W_{p}$ (see e.g. [Vil09, Theorem 6.9]) and note that $\left\|\mu\right\|_{p}=W_{p}(\mu,\delta_{x_{0}})$. We denote for two probability measures $\mu,\nu\in{\mathcal{P}}_{p}(B)$,
$\Gamma_{p}(\mu,\nu)=1+\left\|\mu\right\|_{p}^{p}+\left\|\nu\right\|_{p}^{p}.$
By the triangle inequality, we have $d(y,x_{0})^{p}\leq 2^{p-1}(d(x,x_{0})^{p}+d(x,y)^{p})$ and get for any $\rho\in\Pi(\mu,\nu)$
$\left\|\nu\right\|_{p}^{p}=\int_{B}\int_{B}d(y,x_{0})^{p}\rho(dx,dy)\leq 2^{p-1}\left(\left\|\mu\right\|_{p}^{p}+\int_{B}\int_{B}d(x,y)^{p}\rho(dx,dy)\right).$
Taking the infimum over $\rho$, we get $\left\|\nu\right\|_{p}^{p}\leq 2^{p-1}(\left\|\mu\right\|_{p}^{p}+W_{p}^{p}(\mu,\nu))$ and then
$\Gamma_{p}(\mu,\nu)\leq 1+2^{p}\left\|\mu\right\|_{p}^{p}+2^{p}W_{p}^{p}(\mu,\nu)$ (2.2)
The main object of interest in this section is a family of deterministic operators $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B)$, $s\leq t$, such that $\Theta_{s,s}(\mu)=\mu$ for all $\mu\in\mathcal{P}_{p}(B)$. A simple example is when $B$ is a Banach space and $\Theta_{s,t}(\mu)$ is the distribution of a one step Euler scheme with initial distribution $\mu$ (for further details, see e.g. [PP23] or [BQ24]). General approximation schemes are constructed as follows. We consider an increasing sequence of times $\pi=\{t_{i},i\in\mathbb{N}\}$ with $0\leq t_{i}<t_{i+1}$, and we denote $\gamma_{i}=t_{i}-t_{i-1}\in(0,h)$ for some $h\leq 1$. We then define
$\Theta_{t_{0},t_{i}}^{\pi}(\mu)=\Theta_{t_{i-1},t_{i}}\circ....\circ\Theta_{t_{0},t_{1}}(\mu).$ (2.3)
Following the usual approach of decreasing steps for approximating the invariant measure we also assume that $(\gamma_{n})$ is a nonincreasing sequence for $n\in\mathbb{N}^{\ast}$. If $\Theta_{s,t}(\mu)$ represents the law of a one step Euler scheme with initial distribution $\mu$, then $\Theta_{t_{0},t_{i}}^{\pi}(\mu)$ is the distribution of the Euler scheme on the grid $\pi$ at the final time $t_{i}$ with initial law $\mu$.
Remark 2.1.
Let us note here that in general, for law-dependent dynamics, the approximation $\Theta_{s,t}(\mu)$ cannot be directly simulated and has to be approximated as well. For example, the Euler scheme associated to a McKean-Vlasov SDE involves the current distribution that can be approximated by means of a particle systems or an ergodic time average. We leave the study of the convergence of a fully simulatable scheme in a related general framework for further research.
In the sequel we will use the notation, for $\delta>0$,
$M_{p}(\Theta,\mu,\delta)=\sup_{\pi:t_{i}-t_{i-1}\leq\delta}\sup_{n}\left\|\Theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p}$
with the supremum taken on all the partitions $\pi$ and all $n\in{\mathbb{N}}.$ The proof of the fact that $M_{p}(\Theta,\mu,\delta)$ is finite is not trivial. So, in Subsection 3.2 we give a criterion (of Foster Lyapunov type) which allows to prove it in several concrete examples. More precisely, if this criterion is verified, then (see (3.6))
$M_{p}(\Theta,\mu,\delta)\leq\left\|\mu\right\|_{p}^{p}+C.$ (2.4)
However there are other examples (see the section on the Langevin equation) where this result is already given and consequently we do not need the Foster Lyapunov criterion.
Definition 2.2.
If there exists $\delta>0$ such that $M_{p}(\Theta,\mu,\delta)<\infty$ for every $\mu\in\mathcal{P}_{p}(B)$ we say that $\Theta$ is $p$-bounded.
2.1. Coupling
We now introduce the main assumption on the coupling of two families of operators $\Theta^{1}_{s,t},\Theta^{2}_{s,t}:{\mathcal{P}}_{p}(B)\to{\mathcal{P}}_{p}(B)$, $0\leq s\leq t$.
Definition 2.3.
Let $b_{\ast}>0$ and $\varepsilon>0$. We say that two families of deterministic operators $\Theta^{1}$ and $\Theta^{2}$ are $(p,b_{\ast},\varepsilon)$-coupled if one may find some constants $C_{\ast}>0$ and $h\in(0,1)$ such that, for every $s\leq t\leq s+h$ and every $\mu^{i}\in\mathcal{P}_{p}(B)$, $i=1,2$,
$W_{p}^{p}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))\leq(1-b_{\ast}(t-s))W_{p}^{p}(\mu^{1},\mu^{2})+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}.$ (2.5)
Last, we say that a family $\Theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$, $0\leq s\leq t$, is $(p,b_{\ast},\varepsilon)$-self-coupled if (2.5) holds with $\Theta^{1}=\Theta^{2}=\Theta$.
Remark 2.4.
The really important constants in the above definition are $b_{\ast}$ and $\varepsilon>0$ which appear in (2.5). They determine the speed of convergence in our approximation results, see (2.18) and (2.20). In contrast, the constant $C_{\ast}$ only influences the constants multiplying the speed on convergence in (2.18) and (2.20), and in this sense is less important.
In the sequel we will consider a decreasing sequence $\gamma_{n}\in(0,1),n\in{\mathbb{N}}$ such that $\sum_{n\geq 1}\gamma_{n}=\infty$ and we denote
$\pi=\{t_{n},n\in{\mathbb{N}}\}\quad\text{with}\quad t_{n}=t_{0}+\sum_{i=1}^{n}\gamma_{i}.$ (2.6)
We suppose that
$\overline{\omega}(\pi):=\limsup_{n\rightarrow\infty}\frac{\gamma_{n}-\gamma_{n+1}}{\gamma_{n+1}^{2}}\in(0,\infty).$
For $b>0$ and $\varepsilon>0$ we define
$\sigma_{b,\varepsilon}(n):=\sum_{i=1}^{n}e^{-b(t_{n}-t_{i})}\gamma_{i}^{1+\varepsilon}.$ (2.7)
We need the following technical lemma.
Lemma 2.5.
A. One may find $C_{b,\varepsilon}$ such that for all $n\in\mathbb{N}^{*}$
$\sigma_{b,\varepsilon}(n)$ $\leq$ $C_{b,\varepsilon}\gamma_{n}^{\varepsilon}\quad\text{if}\quad\frac{b}{\varpi(\pi)}>\varepsilon$ (2.8)
$\leq$ $C_{b,\lambda}\gamma_{n}^{\lambda}\quad\text{if}\quad\lambda<\frac{b}{\varpi(\pi)}\leq\varepsilon.$ (2.9)
The constant $C_{b,\varepsilon}$ depends on the sequence $(\gamma_{n})_{n\in{\mathbb{N}}}$ as well, but does not depend on $n$.
B. For the particular case $\gamma_{n}=\frac{1}{n+h^{-1}}$, one has $\varpi(\pi)=1$ and
$\sigma_{b,\varepsilon}(n)\leq\begin{cases}C_{b,\varepsilon}n^{-b\wedge\varepsilon}\text{ if }b\neq\varepsilon,\\ C_{b,\varepsilon}n^{-\varepsilon}\ln(1+n)\text{ if }b=\varepsilon.\end{cases}$
Besides, for $s<t$, $t_{n}=s+\sum_{i=1}^{n}\gamma_{i}$ and $n=n(t)\in{\mathbb{N}}$ such that $t_{n}\leq t<t_{n+1}$, one may find a constant that we still denote $C_{b,\varepsilon}$ such that
$\sigma_{b,\varepsilon}(n)\leq C_{b,\varepsilon}e^{-b\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b=\varepsilon}}).$ (2.10)
A similar result has been used by Pagès and Panloup [PP23, Lemma A.3(i)] with $\varepsilon=1$ and later on by Bally and Qin [BQ24, Lemma A.1(B)]. We give a proof of our slightly different version in the Appendix A.
Lemma 2.6.
Let $\Theta^{1}$ and $\Theta^{2}$ be families which are $(p,b_{\ast},\varepsilon)$-coupled for some $b_{\ast},\varepsilon>0$. We note $C_{\ast}>0$, $h\in(0,1)$ some constants such that (2.5) holds for $\Theta_{s,t}^{1}$ and $\Theta_{s,t}^{2}$, $s\leq t\leq s+h$. We also assume that $\Theta^{1}$ is $p$-bounded and $h\in(0,1)$ is sufficiently small to have $M_{p}(\Theta^{1},\mu,h)<\infty$ and, with $\gamma_{n}=\frac{1}{n+h^{-1}}$,
$2^{p}C_{\ast}\sigma_{b_{\ast},\varepsilon}(k)\leq\frac{1}{2},\quad k\leq n.$ (2.11)
Then, for every $n\in{\mathbb{N}}$, we have
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))$ (2.12)
$\leq e^{-b_{\ast}(t_{n}-t_{0})}W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}(1+M_{p}(\Theta^{1},\mu^{1},h)+W_{p}^{p}(\mu_{1},\mu_{2}))\sigma_{b_{\ast},\varepsilon}(n).$
Besides, $M_{p}(\Theta^{2},\mu,h)<\infty$ for every $\mu\in\mathcal{P}_{p}(B)$ and $\Theta^{2}$ is $p$-bounded.
Proof.
Step 1. We denote $a_{k}=$ $W_{p}^{p}(\Theta_{t_{0},t_{k}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k}}^{2,\pi}(\mu^{2}))$ and we write, for $n\geq 1$,
$a_{n}=e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+\sum_{k=1}^{n}(e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}).$
We define
$M_{n}=\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}\leq M_{p}(\Theta^{1},\mu^{1},h).$
Using the coupling property (2.5) first and (2.2) then, we obtain
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$
$=$ $e^{-b_{\ast}(t_{n}-t_{k})}(1-e^{-b_{\ast}\gamma_{k}})a_{k-1}+e^{-b_{\ast}(t_{n}-t_{k})}(a_{k}-a_{k-1})$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}(b_{\ast}\gamma_{k}a_{k-1}-b_{\ast}\gamma_{k}a_{k-1}+C_{\ast}\Gamma_{p}(\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu^{2}))\gamma_{k}^{1+\varepsilon})$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}+2^{p}W_{p}^{p}(\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu^{2}))\right)\gamma_{k}^{1+\varepsilon}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}M_{n}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum, we get
$a_{n}$ $\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}M_{n}\right)\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}+C_{\ast}2^{p}\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}a_{k-1}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}M_{n}+2^{p}\max_{k\leq n-1}a_{k}\right)\sigma_{b_{\ast},\varepsilon}(n)$
$\leq$ $a_{0}+\frac{1}{2}\left(2^{-p}+M_{n}\right)+\frac{1}{2}\times\max_{k\leq n-1}a_{k}$
the last inequality being a consequence of our hypothesis (2.11). This gives
$\max_{i\leq n}a_{i}\leq a_{0}+\frac{1}{2}\left(1+M_{n}\right)+\frac{1}{2}\times\max_{k\leq n-1}a_{k}$
and finally
$\max_{i\leq n}a_{i}\leq 2a_{0}+M_{n}+1$ (2.14)
Step 2.We come back to (2.1) and we use (2.2) and (2.14) in order to get
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}M_{n}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}2^{p+1}C_{\ast}\left(1+M_{n}+a_{0}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum and using Lemma 2.5, we conclude with
$a_{n}$ $\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+M_{n}+a_{0})\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}$
$=$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+M_{n}+a_{0})\sigma_{b_{\ast},\varepsilon}(n).$
Now, we observe that
$\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu)\right\|_{p}^{p}\leq 2^{p-1}M_{n}+2^{p-1}\max_{k\leq n}a_{k},$
which gives, by (2.14), $M_{p}(\Theta^{2},\mu,h)\leq 2^{p}M_{p}(\Theta^{1},\mu,h)+2^{p}a_{0}+2^{p-1}<\infty$ for every $\mu\in\mathcal{P}_{p}(B)$. ∎
Corollary 2.7.
We suppose that the hypotheses of the previous lemma are in force and we consider the grid $t_{n}=t_{0}+\sum_{i=1}^{n}\frac{1}{i+h^{-1}}.$ Then, for $b_{*}\neq\varepsilon$,
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq$ $An^{-b_{\ast}\wedge\varepsilon},\$(2.15)
with $A=\left((2^{p+1}C_{*}C_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\Theta^{1},\mu^{1},h))\right)$. In the case $b=\varepsilon$, one has to multiply this upper bound by $1+\ln(1+n)$. [^2] In the same way, we have $W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq A(1+(t_{n}-t_{0}))e^{-b_{\ast}\wedge\varepsilon(t_{n}-t_{0})}.$
This is a direct application of Lemma 2.6 and Lemma 2.5 B.
2.2. Invariant measure
Let $\theta_{s,t}$ $\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ be a flow, that is $\theta_{r,t}\circ\theta_{s,r}=\theta_{s,t},s\leq r\leq t$.
Remark 2.8.
For a flow $\theta$, we have
$M_{p}(\theta,\mu,\delta)=\sup_{\pi:t_{i}-t_{i-1}\leq\delta}\sup_{n}\left\|\theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p}=\sup_{t_{0}<t_{n}}\left\|\theta_{t_{0},t_{n}}(\mu)\right\|_{p}^{p}.$
This does not depend on $\delta$, and we simply note this quantity $M_{p}(\theta,\mu)$ in this case.
We look in this subsection for hypotheses which guarantee that $\theta$ admits an invariant probability measure $\nu$, i.e. $\nu=\theta_{s,t}(\nu)$ for every $s<t$. Obviously, if such a $\nu$ exists, it has to be a stationary law since $\nu=\lim_{t\to\infty}\theta_{s,t}(\nu)$. We start with a preliminary estimate.
Corollary 2.9.
Let $\theta_{s,t}$ be a flow which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$-self-coupled for some $b_{\ast},\varepsilon>0$. We note $C_{\ast}>0$, $h\in(0,1)$ some constants such that (2.5) and (2.11) hold. Then, for every $\mu^{1},\mu^{2}\in\mathcal{P}_{p}(B)$ and every $s<t$
$W_{p}^{p}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))\leq Ae^{-b_{\ast}\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b_{*}=\varepsilon}})$
where $A=\left((2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}(1+M_{p}(\theta,\mu^{1}))\right)$ and $\tilde{C}_{b_{\ast},\varepsilon}=C_{b_{\ast},\varepsilon}+1$.
Proof.
Let $s<t$ and $\gamma_{k}=\frac{1}{k+h^{-1}},t_{k}=t_{0}+\gamma_{1}..+\gamma_{k}.$ Let $n\in{\mathbb{N}}$ be such that $t_{n}\leq t<t_{n+1}$. Then, we consider the grid $t_{k}^{\prime}=t_{k}$ for $k\in\{0,\dots,n\}$ and $t_{n+1}^{\prime}=t\leq t_{n+1}$. This means that $\gamma_{k}^{\prime}=\gamma_{k}$ for k=1,,nk=1,...,n and $\gamma_{n+1}^{\prime}<\gamma_{n+1}.$ Recall that, by (2.10), $\sigma_{b_{*},\varepsilon}(n)\leq C_{b_{*},\varepsilon}e^{-b_{*}\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b_{*}=\varepsilon}})$ and by (A.2) $\gamma_{n}\leq h^{-1}e^{-(t_{n}-s).}.$ Then, if $\sigma_{b_{*},\varepsilon}^{\prime}(n)$ is associated to $\gamma_{i}^{\prime},i=0,1,$..., we have
$\sigma_{b_{\ast},\varepsilon}^{\prime}(n+1)$ $=$ $\sum_{i=1}^{n+1}e^{-b_{\ast}(t_{n+1}^{\prime}-t_{0})}(\gamma_{i}^{\prime})^{1+\varepsilon}$
$=$ $e^{-b_{\ast}(t_{n+1}^{\prime}-t_{n})}\sigma_{b_{\ast},\varepsilon}(n)+(\gamma^{\prime}_{n+1})^{1+\varepsilon}$
$\leq$ $(C_{b_{*},\varepsilon}+1)e^{-b_{*}\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b_{*}=\varepsilon}}).$
Since $\gamma_{k}\leq h$ one can thus use Lemma 2.6 to get
$W_{p}^{p}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))=W_{p}^{p}(\theta_{t_{0},t_{n+1}^{\prime}}(\mu^{1}),\theta_{t_{0},t_{n+1}^{\prime}}(\mu^{2}))$
$\ \ \leq e^{-b_{\ast}(t_{n+1}^{\prime}-t_{0})}W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}(1+M_{p}(\theta,\mu^{1})+W_{p}^{p}(\mu_{1},\mu_{2}))\sigma^{\prime}_{b_{\ast},\varepsilon}(n+1).$
Then, the above inequalities give the claim. ∎
Example 2.10.
Let us consider the following time-dependent Ornstein-Uhlenbeck process
$X_{s,t}=X_{s}+\int_{s}^{t}k(\zeta(u)-X_{u})du+\sigma(W_{t}-W_{s}),\ s\leq t,$
with $k,\sigma>0$ and $\zeta:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ being a bounded function. Here, $X_{s}$ denotes a square integrable random variable that is independent of $(W_{t}-W_{s},t\geq s)$. By standard calculations, $X_{s,t}=X_{s}e^{-k(t-s)}+\int_{s}^{t}\zeta(u)e^{-k(t-u)}du+\sigma\int_{s}^{t}e^{-k(t-u)}dW_{u}$. The associated flow $\theta_{s,t}(\mu)=\mathcal{L}(X_{s,t})$ when $X_{s}\sim\mu\in\mathcal{P}_{2}({\mathbb{R}})$ satisfies for $\mu_{1},\mu_{2}\in\mathcal{P}_{2}({\mathbb{R}})$
$W_{2}^{2}(\theta_{s,t}(\mu_{1}),\theta_{s,t}(\mu_{2}))\leq W_{2}^{2}(\mu_{1},\mu_{2})e^{-2k(t-s)}.$
Let $\nu_{s,t}=\mathcal{N}(\int_{s}^{t}\zeta(u)e^{-k(t-u)}du,\frac{\sigma^{2}}{2k})$ be the law of
$\int_{s}^{t}\zeta(u)e^{-k(t-u)}du+\sigma\int_{s}^{t}e^{-k(t-u)}dW_{u}+\sqrt{\frac{\sigma^{2}}{2k}e^{-2k(t-s)}}Y$
with $Y\sim\mathcal{N}(0,1)$ independent of the other variables. We have
$W^{2}_{2}(\theta_{s,t}(\mu),\nu_{s,t})\leq 2\mathbb{E}[X_{s}^{2}]e^{-2k(t-s)}+\frac{\sigma^{2}}{k}e^{-2k(t-s)}=_{t\to\infty}O(e^{-2k(t-s)}).$
However, we can choose $\zeta$ so that $\int_{s}^{t}\zeta(u)e^{-k(t-u)}du$ does not converge as $t\to\infty$ (take for example $\zeta(u)=\cos(u)$) and therefore $\nu_{s,t}$ (and thus $\theta_{s,t}(\mu)$) does not weakly converge.
As illustrated by this simple example, time-inhomogenous flow may not have a stationary distribution even though they are self-coupled. To go further, we consider time homogeneous flows which means that
$\theta_{s,t}(\mu)=\theta_{0,t-s}(\mu)\text{ for }0\leq s\leq t.$
From this, we get easily $\theta_{0,t}\circ\theta_{0,t^{\prime}}=\theta_{0,t+t^{\prime}}$ for any $t,t^{\prime}\geq 0$ and thus
$\theta_{t,t+\delta}\circ\theta_{s,t}=\theta_{0,t-s+\delta}=\theta_{s,t}\circ\theta_{t,t+\delta},\ \delta\geq 0,\ t\geq s\geq 0.$ (2.16)
In addition, we consider a flow $\theta_{s,t}$ which is continuous for $W_{p}$, in the sense that for every fixed $s\leq t$
$\lim_{n\rightarrow\infty}W_{p}(\mu_{n},\mu)=0\quad\Rightarrow\quad\lim_{n\rightarrow\infty}W_{p}(\theta_{s,t}(\mu_{n}),\theta_{s,t}(\mu))=0.$ (2.17)
Proposition 2.11.
Consider $\theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$ a continuous time-homogeneous flow, which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0.$ Then, there exists a unique probability measure $\nu\in\mathcal{P}_{p}(B)$ such that for every $s\leq t$ one has $\nu=\theta_{s,t}(\nu).$ It satisfies $\|\nu\|_{p}^{p}=M_{p}(\theta,\nu)$ and we have for every $s\leq t$ and every $\mu\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (2.18)
with $A$ being as given by Corollary 2.9.
Proof.
Let $s\leq t,\delta>0$ and $\mu\in\mathcal{P}_{p}(B)$. We have
$W_{p}(\mu,\theta_{t,t+\delta}(\mu))\leq\left\|\mu\right\|_{p}+\left\|\theta_{t,t+\delta}(\mu)\right\|_{p}\leq 2M_{p}(\theta,\mu)$
so that, using the time homogeneous flow property (2.16) and Corollary 2.9 with $\mu^{1}=\mu$ and $\mu^{2}=\theta_{t,t+\delta}(\mu)$
$W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t+\delta}(\mu))=W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t}(\theta_{t,t+\delta}(\mu)))\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s))^{\mathbf{1}_{b_{\ast}=\varepsilon}}.$ (2.19)
Therefore, if $t_{n}\rightarrow\infty$ , we get that $\theta_{s,t_{n}}(\mu)$, $n\in{\mathbb{N}}$ is a Cauchy sequence. Since $\mathcal{P}_{p}(B)$ is complete with respect to $W_{p}$, we define its limit in the Wasserstein metric as $\nu:=\lim_{n\rightarrow\infty}\theta_{s,t_{n}}(\mu)\in\mathcal{P}_{p}(B)$. Coming back to 2.19 and passing to the limit we get
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s))^{\mathbf{1}_{b_{\ast}=\varepsilon}},$
This proves (2.18).
We now check that $\nu=\lim_{n\rightarrow\infty}\theta_{s^{\prime},t_{n}}(\mu)$ for any $s^{\prime}\geq 0$. For $s^{\prime}\leq s$, we write for $n$ sufficiently large
$W_{p}^{p}(\theta_{s^{\prime},t_{n}}(\mu),\theta_{s,t_{n}}(\mu))=W_{p}^{p}(\theta_{s,t_{n}}(\theta_{s^{\prime},s}(\mu)),\theta_{s,t_{n}}(\mu))\rightarrow 0,$
by using Corollary 2.9. We proceed similarly for $s^{\prime}>s$.
This property gives easily that $\nu$ is invariant: we get indeed from the continuity property (2.17) that
$\theta_{s,t}(\nu)=\lim_{n\to\infty}\theta_{s,t}(\theta_{t,t_{n}}(\mu))=\lim_{n\to\infty}\theta_{s,t_{n}}(\mu)=\nu.$
Let us prove uniqueness. Let $\nu^{\prime}\in\mathcal{P}_{p}(B)$ be another invariant measure. Then, using Corollary 2.9, $W^{p}_{p}(\nu,\nu^{\prime})=W_{p}^{p}(\theta_{s,t}(\nu),\theta_{s,t}(\nu^{\prime}))\rightarrow 0$ as $t\rightarrow\infty$. ∎
We discuss now the rate of convergence to the invariant measure using an iterative scheme algorithm.
Theorem 2.12.
Let $\theta$ be a time homogeneous continuous flow on ${\mathcal{P}}_{p}(B)$ which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0$ and $\nu$ the invariant measure obtained in Proposition 2.11. Let $\Theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$, $s\leq t$, be a family of applications which is $(p,b_{\ast},\varepsilon)$ coupled with $\theta_{s,t}$. Let $h\in(0,1)$ be such that both (2.11) and the coupling property (2.5) between $\theta$ and $\Theta$ hold. Let $\gamma_{n}=\frac{1}{n+h^{-1}}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, we have for every $\mu\in\mathcal{P}_{p}(B)$
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq An^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (2.20)
with $A=2^{2p}C_{*}(1+M_{p}(\theta,\mu))(2C_{b_{\ast},\varepsilon}+1)$.
Proof.
We suppose that $b_{*}\neq\varepsilon$: if not, we multiply the upper bound with $1+\ln(1+n)$. By (2.15), we have
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\theta_{s,t_{n}}(\mu))\leq A^{\prime}n^{-b_{\ast}\wedge\varepsilon},$
with $A^{\prime}=2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\theta,\mu))$. By (2.18), we also have
$W_{p}^{p}(\theta_{s,t_{n}}(\mu),\nu)\leq A^{\prime\prime}e^{-\varepsilon\wedge b_{\ast}(t_{n}-s)}\leq A^{\prime\prime}n^{-\varepsilon\wedge b_{\ast}}$
with $A^{\prime\prime}=2^{p+1}C_{*}(C_{b_{\ast},\varepsilon}+1)(1+M_{p}(\theta,\mu))$, since $t_{n}-s\leq\ln(1+n)$. Using the triangular inequality and $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$, we get the claim. ∎
Remark 2.13.
Now that we have stated in Theorem 2.12 the main result related to our abstract framework, it is interesting to compare with the recent paper of Schuh and Souttar [SS24] that came out while working on this paper. As [SS24], our goal is to give a general framework to study the approximation error of a stationary measure. However, there are significant differences between our works, and we list here the main ones.
  1. (1)
    The framework in [SS24] considers approximations defined for all $t\geq 0$ and parametrized by a parameter $\delta>0$. When applied to Euler type approximation schemes (see [SS24, Subsection 3.2]), this parameter represents the constant time step. In contrast, our framework is meant for approximation schemes and a decreasing sequence of time steps.
  2. (2)
    [SS24, General assumptions 1 (1)] asks for contractivity, which implies the exponential convergence to the invariant law of one process. In our setting, the self-coupling property (2.5) on $\theta$ implies this property in Proposition 2.11. Note that [SS24, General assumptions 1 (1)] implies condition (2.5), but the reciprocal is not true: $W_{p}^{p}(\mu^{1},\mu^{2})$ is trivially upper bounded by $\Gamma_{p}(\mu^{1},\mu^{2})$, but we cannot upper bound $\Gamma_{p}(\mu^{1},\mu^{2})$ by $CW_{p}^{p}(\mu^{1},\mu^{2})$. Condition (2.5) allows the error term in $(t-s)^{1+\varepsilon}$ which is convenient in order to deal with Euler type schemes, see in particular the example concerning the Boltzmann type equation in Subsection 5.3 below.
  3. (3)
    [SS24, General assumptions 1 (2)] can seen as a consequence of the coupling assumption (2.5) between $\theta$ and $\Theta$, provided that $\theta$ and $\Theta$ are $p$-bounded. Indeed, if we take $\delta<h$ and set $a_{k}=W_{p}^{p}(\theta^{\pi}_{s,s+k\delta}(\mu),\Theta^{\pi}_{s,s+k\delta}(\mu))$ with $\pi=\{s+n\delta,n\in{\mathbb{N}}\}$ and $\mu\in\mathcal{P}_{p}(B)$, we get $a_{0}=0$ and $a_{k+1}\leq e^{-b_{*}\delta}a_{k}+C\delta^{1+\varepsilon}$. This gives $a_{k}\leq C\frac{\delta^{1+\varepsilon}}{1-e^{-b_{*}\delta}}$ for all $k\in{\mathbb{N}}$, and thus $W_{p}(\theta^{\pi}_{s,s+k\delta}(\mu),\Theta^{\pi}_{s,s+k\delta}(\mu))\leq(C/b_{*})^{1/p}\delta^{\varepsilon/p}$ when $\delta$ is smaller than $1/b_{*}$. Note also that we directly get an estimate for all $k\geq 0$ as in [SS24, Theorem 1.1].
    Let us stress also that [SS24, General assumptions 1 (2)] involves the multi step approximation scheme. Instead, our error hypothesis (2.5) only concerns the one step approximation scheme, which is crucial then in order to deal with decreasing time steps. In practice, it is often easier to check one step estimates : though [SS24, General assumptions 1 (2)] is weaker, one would typically use one step estimates and adapt the arguments above to prove it.
Thus, modulo $p$-boundedness assumptions, [SS24, General assumptions 1 (1)] is a bit stronger than the self-coupling property on $\theta$ while [SS24, General assumptions 1 (2)] is a bit weaker than our coupling assumption between $\theta$ and $\Theta$. Our assumptions are thus different from those of [SS24]. They allow us to consider a decreasing sequence of time steps and then to get a rate of convergence, which does not seem possible with [SS24, General assumptions 1]. Note that considering a decreasing sequence of time steps is somehow related to the question raised in [SS24, Remark 2.5] on the distribution limit when both $t\to\infty$ and $\delta\to 0$ and the commutativity of the limits.
2.3. Probabilistic representation
Now, we give a probabilistic framework which allows to prove that $\Theta^{1}$ and $\Theta^{2}$ are $(p,b_{\ast},\varepsilon)$- coupled (and which will be used in the sequel in our examples).
Definition 2.14.
We say that we have a “joint probabilistic representation” for $(\Theta^{1},\Theta^{2})$ if the following holds. There exists some $h\in(0,1)$ such that for every $s<t$ with $t-s\leq h$ and $\mu^{1},\mu^{2}\in\mathcal{P}_{p}(B)$ one may find a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ (which may depend on $s<t$ and $\mu^{1},\mu^{2})$ and some $B$-valued $p$-integrable random variables $X^{1},X^{2}$ and $\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2}$ such that
$X^{i}\sim\mu^{i}\quad\text{and}\quad\mathcal{X}_{s,t}^{i}\sim\Theta_{s,t}^{i}(\mu^{i})\quad i=1,2$
and the following two conditions are satisfied for some positive constants $\varepsilon$, $b_{\ast}$, $C_{\ast}$, and any $0<s\leq t$ with $t-s\leq h$
$W_{p}^{p}(\mu^{1},\mu^{2})=$ $\mathbb{E}(d^{p}(X^{1},X^{2})),$ (2.21)
$\mathbb{E}(d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2}))\leq$ $(1-b_{\ast}(t-s))\mathbb{E}(d^{p}(X^{1},X^{2}))+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon},$ (2.22)
When $\Theta^{1}=\Theta^{2}$, we say that we have a self joint probabilistic representation for $\Theta^{1}$, which echoes to Definition 2.3 where coupling and self coupling are introduced.
We refer to Villani [Vil09, Theorem 4.1] for the existence of an optimal coupling, so that condition (2.21) can always be achieved on an atomless probability space.
Proposition 2.15.
Suppose that we have a joint probabilistic representation for $\Theta^{1}$ and $\Theta^{2}$. Then, $\Theta^{1}$ and $\Theta^{2}$ are $(p,b_{\ast},\varepsilon)$-coupled .
Proof.
We need to check the conditions in Definition 2.3. Using (2.22) first and (2.21) we have
$W_{p}^{p}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))$ $\leq$ $\mathbb{E}(d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2}))$
$\leq$ $(1-b_{\ast}(t-s))\mathbb{E}(d^{p}(X^{1},X^{2}))+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}$
$=$ $(1-b_{\ast}(t-s))W_{p}^{p}(\mu^{1},\mu^{2})+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}$
so (2.5) is proved. ∎
Now, we assume that $\Theta$ has a probabilistic representation as in Definition 2.14:  for every $s<t$ and every $\mu\in\mathcal{P}_{p}(B),$ one may find a probability space $(\Omega,\mathcal{F},{\mathbb{P}})$ (which may depend on $s<t$ and $\mu)$ and some $B$ valued square integrable random variables $X$ and $\mathcal{X}_{s,t},$ such that $X\sim\mu$ and $\mathcal{X}_{s,t}\sim\Theta_{s,t}(\mu).$ In particular
$\left\|\mathcal{X}_{s,t}\right\|_{p}=\left\|\Theta_{s,t}(\mu)\right\|_{p}\quad\left\|X\right\|_{p}=\left\|\mu\right\|_{p}$
Lemma 2.16.
Let $h>0$ and suppose that $\Theta$ has a probabilistic representation satisfying
$\left\|\mathcal{X}_{s,t}\right\|_{p}^{p}\leq(1-b(t-s))\left\|X\right\|_{p}^{p}+C(t-s)$
for some $b$ and $C$ and any $t-s\in[0,h]$. Then $\Theta$ verifies the $p$-Foster-Lyapunov condition (3.2) and consequently is $p$-bounded, and by (3.6)
$M_{p}(\Theta,\mu)\leq\left\|\mu\right\|_{p}^{p}+C\frac{e^{b}}{b}.$ (2.23)
3. Some useful tools
3.1. Using the sewing lemma to prove coupling
In this section, we explain how the sewing lemma can be used to get automatically some assumptions that are required in Theorem 2.12 when working with the $W_{1}$ distance. We start by the following motivating lemma that requires to work with $p=1$, as explained after the proof.
Lemma 3.1.
Let $\Theta^{i}_{s,t}:{\mathcal{P}}_{1}(B)\rightarrow{\mathcal{P}}_{1}(B)$, $s\leq t$, $i=1,2$ be two families of applications such that:
  • $\Theta^{1}$ is $1$-bounded and $(1,b_{\ast},\varepsilon)$ self-coupled,
  • There exists $C\in{\mathbb{R}}_{+}$ and $h>0$ such that for all $0\leq t-s\leq h$ and $\mu\in\mathcal{P}_{1}(B)$, $W_{1}(\Theta^{1}_{s,t}(\mu),\Theta^{2}_{s,t}(\mu))\leq C(1+\|\mu\|_{1})(t-s)^{1+\varepsilon}$.
Then, $\Theta^{1}$ and $\Theta^{2}$ are $(1,b_{\ast},\varepsilon)$-coupled. Furthermore, $\Theta^{2}$ is $1$-bounded and $(1,b_{\ast},\varepsilon)$ self-coupled as well.
Proof.
Let $\mu^{1},\mu^{2}\in\mathcal{P}_{1}(B)$. We use the triangle inequality, the fact that $\Theta^{1}$ is $(1,b_{\ast},\varepsilon)$ self-coupled and the assumption to get
$W_{1}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))$ $\leq W_{1}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{1}(\mu^{2}))+W_{1}(\Theta_{s,t}^{1}(\mu^{2}),\Theta_{s,t}^{2}(\mu^{2}))$
$\leq(1-b_{*}(t-s))W_{1}(\mu^{1},\mu^{2})+C_{*}\Gamma_{1}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}$
$\quad+C(1+\|\mu^{2}\|_{1})(t-s)^{1+\varepsilon}.$
We thus get that $\Theta^{1}$ and $\Theta^{2}$ are $(1,b_{\ast},\varepsilon)$-coupled. And by Lemma 2.6, $\Theta^{2}$ is also $1$-bounded. ∎
We see here why we need $p=1$. For $p>1$, we would apply the triangular inequality and then $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{b})$: the multiplicative factor $2^{p-1}>1$ prevents to get the coupling property (2.5).
We now present the sewing lemma. To do so, we define
$\mathcal{E}_{1}(\mathcal{P}_{1}(B))=\left\{\Theta:\mathcal{P}_{1}(B)\rightarrow\mathcal{P}_{1}(B):\sup_{\mu\in\mathcal{P}_{1}(B)}\frac{\|\Theta(\mu)\|_{1}}{1+\|\mu\|_{1}}<\infty\right\},$
that we endow with the following distance
$\mathcal{D}_{1}(\Theta,\overline{\Theta})=\sup_{\mu\in\mathcal{P}_{1}(B)}\frac{W_{1}(\Theta(\mu),\overline{\Theta}(\mu))}{1+\|\mu\|_{1}}.$
From [Vil09, Theorem 6.18], $(\mathcal{P}_{1}(B),W_{1})$ is a complete metric space. By [AB23, Lemma B.1], $(\mathcal{E}_{1}(\mathcal{P}_{1}(B)),\mathcal{D}_{1})$ is a complete metric space. For $U,V\in\mathcal{E}_{1}(\mathcal{P}_{1}(B))$, we note $UV=U\circ V$ and we check easily that $UV\in\mathcal{E}_{1}$. From [AB23, Lemma 2.1], we have the following sewing lemma.
Lemma 3.2.
(Sewing lemma) Let $T>0$ and suppose that $(\Theta_{s,t})_{0\leq s\leq t\leq T}$ is a family such that $\Theta_{s,t}\in\mathcal{E}_{1}$ and:
  1. (1)
    $\sup_{0\leq s\leq t\leq T}\mathcal{D}_{1}(\Theta_{s,t},Id)<\infty$,
  2. (2)
    there exists $C_{\textup{lip}}$ such that for any $0\leq s\leq t\leq T$, $n\in\mathbb{N}^{\ast}$, any time grid $\pi=\left\{t_{0}=s<t_{1}<\dots<t_{n}=t\right\}$ and every $U,\tilde{U}\in\mathcal{E}_{1}(\mathcal{P}_{1}(B))$,
[…]
  3. (3)
    there exists $C_{\textup{sew}}$ and $\varepsilon>0$ such that for any $0\leq s<u<t<T$ ,
    $\mathcal{D}_{1}(\Theta_{s,t}U,\Theta_{u,t}\Theta_{s,u}\tilde{U})\leq C_{\textup{sew}}(t-s)^{1+\varepsilon}.$
Then, there exists a unique family $\theta_{s,t}\in\mathcal{E}_{1}(\mathcal{P}_{1}(B))$ which is a flow such that
$\exists C>0,\ \forall s\leq t\leq T,\ \mathcal{D}_{1}(\theta_{s,t},\Theta_{s,t})\leq C(t-s)^{1+\varepsilon}.$
[…]
Therefore, the sewing lemma gives the estimate
$W_{1}(\theta_{s,t}(\mu),\Theta_{s,t}(\mu))\leq C(t-s)^{1+\varepsilon}(1+\|\mu\|_{1})$ (3.1)
that is needed to apply Lemma 3.1. More precisely, Lemma 3.1 can be then applied in two ways:
  • If $\Theta$ is $1$-bounded and $(1,b_{\ast},\varepsilon)$ self-coupled, then $\Theta$ and $\theta$ are $(1,b_{\ast},\varepsilon)$-coupled and $\theta$ is also $1$-bounded.
  • If we are able to prove independently (relying on continuous time arguments) that $\theta$ is $1$-bounded and is $(1,b_{\ast},\varepsilon)$ self-coupled, then $\Theta$ and $\theta$ are $(1,b_{\ast},\varepsilon)$-coupled and $\Theta$ is $1$-bounded. In particular, if $\theta$ is time homogeneous and $\nu$ is the invariant measure for $\theta$, and if we take the time sequence $t_{k}=s+\sum_{i=1}^{k}\frac{1}{i+h^{-1}}$ with $h\in(0,1)$ as in Theorem 2.12 then, by (2.20), for every $\mu\in\mathcal{P}_{1}(B)$,
    $W_{1}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq An^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}}).$
3.2. A Foster-Lyapunov type criterion to prove $p$-boundedness
We consider now a family of operators $\Theta_{s,t}:\mathcal{P}_{p}(B)\to\mathcal{P}_{p}(B)$ and we discuss the boundedness property. For a $B$-valued random variable $X$, we note $\|X\|_{p}=\mathbb{E}[d(X,x_{0})^{p}]^{1/p}$.
Definition 3.3.
We say that a family of deterministic operators $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B)$, satisfies the $p$-Foster-Lyapunov condition if there exist some constants $\overline{b},\overline{C},h\in(0,1)$, such that for every $s\leq t\leq s+h$ and every $\mu\in\mathcal{P}_{p}(B)$,
$\left\|\Theta_{s,t}(\mu)\right\|_{p}^{p}\leq(1-\overline{b}(t-s))\left\|\mu\right\|_{p}^{p}+\overline{C}(t-s).$ (3.2)
Note that if we prove (3.2) for some distance $d(x,y)$, this does not imply that the same property is satisfied for an equivalent distance due to the factor $(1-\overline{b}(t-s))$ in the property (3.2). Thus, in some cases, one may be able to prove (3.2) for some special distance $d$, but not for $|\cdot|$, or conversely. The constants throughout the article depend on $p$ and $d$ without making this dependence explicit as we will always consider $p$ and $d$ to be fixed.
If (3.2) holds, then we clearly have
$\left\|\Theta_{s,t}(\mu)\right\|_{p}^{p}\leq e^{-{\bar{b}}(t-s)}\left\|\mu\right\|_{p}^{p}+\overline{C}(t-s).$ (3.3)
Conversely, if (3.3) holds, we get by a Taylor expansion of the exponential function that there exists $0<h^{\prime}<h$ such that for $0\leq t-s\leq h^{\prime}$, (3.2) holds with $\bar{b}/2$ instead of $\bar{b}$. Thus, the conditions (3.2) and (3.3) are essentially equivalent, but we prefer to use (3.2) because we usually directly get this type of estimates on practical examples.
Remark 3.4.
The hypothesis (3.2) is a variant of the Foster-Lyapunov drift condition (see e.g. [MT93], [JH01, Eq. (13)] and [DM17, Eq. (6)]) which is used in the Monte Carlo Markov Chain (MCMC) literature. Let us consider the case of a time homogeneous family, i.e. $\Theta_{s,t}(\mu)=\mu P_{t-s}$, where $P_{\gamma}(x,dy)$, $\gamma>0$ is a probability kernel and $\mu P_{\gamma}(dy)=\int_{B}P_{\gamma}(x,dy)\mu(dx)$ as usual. Following [DM17], a function $V$ satisfy a Foster-Lyapunov drift condition for $P_{\gamma}$ if there exists $\bar{\gamma}>0$, $\lambda\in(0,1)$ and $c\geq 0$ such that $P_{\gamma}V(x)\leq\lambda^{\gamma}V(x)+c\gamma$ for all $\gamma\in(0,\bar{\gamma}]$. If $V(x)=|x|^{p}$ satisfies this Foster-Lyapunov drift condition, then integrating this inequality with respect to $\mu$ gives precisely (3.3) with $\bar{b}=-\log(\lambda)$, $h=\bar{\gamma}$ and d(x,y)=|xy|d(x,y)=|x-y|. Conversely, if (3.3) holds with d(x,y)=|xy|d(x,y)=|x-y| then we get by taking $\mu=\delta_{x}$ that $P_{\gamma}V(x)\leq\lambda^{\gamma}V(x)+c\gamma$ for $V(x)=|x|^{p}$.
Remark 3.5.
Suppose that one may find constants $\widehat{b},\widetilde{C},\widehat{C},{\varepsilon},h>0$ such that for every $s\leq t\leq t+h$ and $\mu\in\mathcal{P}_{p}(B)$,
$\left\|\Theta_{s,t}(\mu)\right\|_{p}^{p}\leq(1-\widehat{b}(t-s))\left\|\mu\right\|_{p}^{p}+\widetilde{C}(t-s)+\widehat{C}(1+\left\|\mu\right\|_{p}^{p})(t-s)^{1+{\varepsilon}}.$ (3.4)
Then, for every $0\leq t-s\leq\min\left(h,\left(\frac{\widehat{b}}{2\widehat{C}}\right)^{1/{\varepsilon}}\right)$, (3.2) holds with $\overline{b}=\widehat{b}/2$, and $\overline{C}=\widetilde{C}+\widehat{C}h^{{\varepsilon}}.$
We consider a family of deterministic operators $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ and recall the notations and definitions given in Section 2. We now prove an estimate which is an iterated version of (3.2).
Lemma 3.6.
Suppose that the family $\Theta_{s,t}$, $s\leq t$ satisfies the $p$-Foster-Lyapunov condition, and let $\overline{b},\overline{C},h>0$ be such that (3.2) holds. Let $(t_{i})_{i\in{\mathbb{N}}}$ be an increasing sequence such that $t_{0}\geq 0$, $t_{1}-t_{0}\leq h$ and $(t_{i}-t_{i-1})_{i\in{\mathbb{N}}^{\ast}}$ is nonincreasing. Then, for every $n\in{\mathbb{N}}$
$\left\|\Theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p}\leq e^{-\overline{b}(t_{n}-t_{0})}\left\|\mu\right\|_{p}^{p}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$ (3.5)
In particular $\Theta$ is $p$-bounded and
$M_{p}(\Theta,\mu)\leq\left\|\mu\right\|_{p}^{p}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$ (3.6)
In addition, assume that $\Theta$ is a flow (i.e. $\Theta_{s,t}=\Theta_{r,t}\circ\Theta_{s,r}$ for every $s<r<t)$. Then, for every $s<t$ and every $\mu\in\mathcal{P}_{p}(B)$, one has
$\left\|\Theta_{s,t}(\mu)\right\|_{p}^{p}\leq e^{-\overline{b}(t-s)}\left\|\mu\right\|_{p}^{p}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$ (3.7)
Proof.
We denote $a_{k}=\left\|\Theta_{t_{0},t_{k}}^{\pi}(\mu)\right\|_{p}^{p}$ and write
$e^{\overline{b}t_{n}}a_{n}=e^{\overline{b}t_{0}}a_{0}+\sum_{k=1}^{n}(e^{\overline{b}t_{k}}a_{k}-e^{\overline{b}t_{k-1}}a_{k-1}).$
Using the property (3.2) (with $t_{k}-t_{k-1}=\gamma_{k}\leq\gamma_{1}\leq h)$
$e^{\overline{b}t_{k}}a_{k}-e^{\overline{b}t_{k-1}}a_{k-1}=e^{\overline{b}t_{k}}(1-e^{-\overline{b}\gamma_{k}})a_{k-1}+e^{\overline{b}t_{k}}(a_{k}-a_{k-1})$
$\leq$ $e^{\overline{b}t_{k}}(\overline{b}\gamma_{k}a_{k-1}-\overline{b}\gamma_{k}a_{k-1}+\overline{C}\gamma_{k})=\overline{C}e^{\overline{b}t_{k}}\gamma_{k}.$
Therefore, we get
$e^{\overline{b}t_{n}}a_{n}\leq e^{\overline{b}t_{0}}a_{0}+\overline{C}\sum_{k=1}^{n}e^{\overline{b}t_{k}}\gamma_{k},$
which gives
$a_{n}$ $\leq e^{-\overline{b}(t_{n}-t_{0})}a_{0}+\overline{C}\sum_{k=1}^{n}e^{-\overline{b}(t_{n}-t_{k})}\gamma_{k}$
$\leq e^{-\overline{b}(t_{n}-t_{0})}a_{0}+\overline{C}e^{\overline{b}}\sum_{k=1}^{n}e^{-\overline{b}(t_{n}-t_{k-1})}\gamma_{k}\leq e^{-\overline{b}(t_{n}-t_{0})}a_{0}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$
The last inequality uses that $e^{\overline{b}t_{k-1}}\gamma_{k}\leq\int_{t_{k-1}}^{t_{k}}e^{\overline{b}s}ds$ and exact integration. So (3.5) and (3.6) are proved.
We now assume that $\Theta$ is a flow. We take $n$ such that (ts)/n<h(t-s)/n<h and consider the uniform grid $t_{k}=s+k\frac{t-s}{n}$ so that $t_{0}=s$ and $t_{n}=t$. We use the previous result for the time grid $\pi=\{t_{k},k\in{\mathbb{N}}\}$ and the flow property $\Theta_{t_{0},t_{n}}^{\pi}=\Theta_{s,t}$ so that (3.7) coincides with (3.5). ∎
In many situations to follow in the examples, one applies a sequence of arguments which are a recall of the results of Theorems 2.15, 2.12 and Proposition 2.11.
Theorem 3.7.
Let $\Theta,\theta:\mathcal{P}_{p}(B)\to\mathcal{P}_{p}(B)$ and that $\theta$ is a time homogeneous continuous flow. Assume that we have two joint probabilistic representations for $(\Theta,\theta)$ and $(\theta,\theta)$ in the sense of Definition 2.14. Furthermore, we assume that they satisfy (3.4). Then the invariant measure associated to $\theta$, denoted by $\nu$ exists and is unique. It satisfies $\|\nu\|_{p}^{p}=M_{p}(\theta,\nu)$ and we have for every $s\leq t$ and every $\mu\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (3.8)
with $A$ being as given by Corollary 2.9.
Furthermore, let $\gamma_{n}=\frac{1}{n+h^{-1}}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, we have for every $\mu\in\mathcal{P}_{p}(B)$
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq An^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (3.9)
with $A=2^{2p}C_{*}(1+M_{p}(\theta,\mu))(2C_{b_{\ast},\varepsilon}+1)$.
4. Examples with non global contraction conditions (Langevin type equations)
The present section is devoted to prove that, once the convergence of the semigroup to the invariant measure is obtained, the estimate of the error done by using Euler scheme, is possible (see Lemma 3.1). We illustrate this by two non trivial examples.
In several recent papers (see Eberle [Ebe16], Schuh [Sch24] ) authors have used a special coupling procedure, called “reflecting coupling” in order to prove convergence to equilibrium for the Langevin equation. An important point is that the result holds under a weak contraction hypothesis (i.e. out of a compact set), in contrast with older results in which uniform construction properties are needed. It turns out that we are able to use Eberle’s result concerning the flow in order to prove convergence of the Euler scheme as well and to control the speed of convergence. First, we consider the classical Langevin equation given by
$dX_{t}=b(X_{t})dt+dB_{t}$ (4.1)
where $B$ is an $m$ dimensional Brownian motion and $b:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ is a Lipschitz continuous function. We assume that
$\left|b(x)-b(y)\right|\leq L_{b}\left|x-y\right|$ (4.2)
and that there exists $R,\kappa>0$ such that
$\left\langle x-y,b(x)-b(y)\right\rangle\leq-\kappa\left|x-y\right|^{2}\quad\text{if}\quad\left|x-y\right|\geq R.$ (4.3)
Then, we are precisely in the framework of [Ebe16, Example 1]. In particular, in [Ebe16, Equation (10)], the author finds a function $f:{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}_{+}$ which is concave, increasing, such that $f(0)=0$, $f^{\prime}(0)=1$ and $0<c_{\ast}\leq f^{\prime}(r)\leq 1$ for every $r$. In what follows, the precise expression of $f$ is not important for us. To the function $f$ one associates the distance
$d_{f}(x,y)=f(\left|x-y\right|)$
which is equivalent with the Euclidean distance: $c_{*}|x-y|\leq d_{f}(x,y)\leq|x-y|$. We note for $\mu,\nu\in\mathcal{P}_{1}({\mathbb{R}}^{d})$
$W_{1}(\mu,\nu)=\inf_{\rho\in\Pi(\mu,\nu)}\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}d_{f}(x,y)\rho(dx,dy),$
the corresponding $1$-Wasserstein distance, and $W^{|\cdot|}_{1}(\mu,\nu)=\inf_{\rho\in\Pi(\mu,\nu)}\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}|x-y|\rho(dx,dy)$ the classical 1-Wasserstein distance. We then have
$c_{*}W^{|\cdot|}_{1}(\mu,\nu)\leq W_{1}(\mu,\nu)\leq W_{1}^{|\cdot|}(\mu,\nu).$ (4.4)
It is equivalent to the standard $1$-Wasserstein distance Then, [Ebe16, Corollary 2 (13) and Corollary 3] gives
Theorem 4.1.
Suppose that (4.2) and (4.3) hold and let $\theta_{s,t}(\mu)$ be the law of $X_{t}$ when $X_{s}\sim\mu$ is independent of $(B_{u}-B_{s},u\geq s)$. Then, there is a constant $b_{*}>0$ such that
$W_{1}(\theta_{s,t}(\mu),\theta_{s,t}(\nu))\leq W_{1}(\mu,\nu)e^{-b_{\ast}(t-s)}.$ (4.5)
Moreover, $\theta$ has a unique invariant measure $\nu$ which has finite moment of order one.
Remark 4.2.
Using that $\theta_{s,t}(\nu)=\nu$ we obtain that $\theta$ is $1$-bounded. Indeed, following (2.1), we set $\left\|\mu\right\|_{1}=\int_{{\mathbb{R}}^{d}}d_{f}(x,0)\mu(dx)$ and for any $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, we have:
$\left\|\theta_{s,t}(\mu)\right\|_{1}$ $\leq$ $\left\|\theta_{s,t}(\nu)\right\|_{1}+W_{1}(\theta_{s,t}(\mu),\theta_{s,t}(\nu))$
$\leq$ $\left\|\nu\right\|_{1}+W_{1}(\mu,\nu),$
by using (4.5).
Remark 4.3.
From (4.4), we get easily
$W_{1}^{|\cdot|}(\theta_{s,t}(\mu),\theta_{s,t}(\nu))\leq\frac{1}{c_{*}}e^{-b_{*}(t-s)}W_{1}(\mu,\nu)$
but then the multiplicative constant becomes $1/c_{*}$ which is in general strictly greater that one (otherwise the distances $W_{1}$ and $W_{1}^{|\cdot|}$ would be the same). This breaks the contraction property that is crucial in the convergence analysis toward the stationary measure. In several recent papers and prominently Eberle [Ebe16], the authors succeed in constructing a distance $d$ that allows to obtain strict contraction with respect to the Wasserstein distance associated to this distance. This is why we include in our framework of Section 2 a general distance $d$.
We now construct the following Euler scheme:
$Y_{s,t}$ $=$ $Y+b(Y)(t-s)+B_{t}-B_{s}$
$\Theta_{s,t}(\mathcal{L}(Y))$ $=$ $\mathcal{L}(Y_{s,t}).$
Lemma 4.4.
Suppose that $b$ satisfies (4.2). Then, there exists $C\in{\mathbb{R}}_{+}$ such that for all $s\leq t\leq s+1$ and $\mu\in\mathcal{P}_{1}({\mathbb{R}}^{d})$,
$W_{1}(\theta_{s,t}(\mu),\Theta_{s,t}(\mu))\leq C(1+\|\mu\|_{1})(t-s)^{3/2}.$ (4.6)
Thus, from Theorem 4.1 we have that Lemma 3.1 holds with $\varepsilon=\frac{1}{2}.$
Proof.
Let $X_{s,t}\equiv X_{s,t}(Y)$ be the solution of
$X_{s,t}=Y+\int_{s}^{t}b(X_{s,r})dr+B_{t}-B_{s},\ t\geq s.$
Then, for $|t-s|\leq 1$, we have $\mathbb{E}[\left|X_{s,t}-Y\right|]\leq C(1+\mathbb{E}\left|Y\right|)(t-s)^{1/2}$, for a constant $C$ depending on $b(0)$ and $L_{b}$. We now combine the Lipschitz property and the previous inequality to get
$\mathbb{E}\left[\left|X_{s,t}-Y_{s,t}\right|\right]\leq\int_{s}^{t}C(1+\mathbb{E}\left|Y\right|)(r-s)^{1/2}dr\leq C(1+\mathbb{E}\left|Y\right|)(t-s)^{3/2}.$
With the above Lemma we can give the error estimate for the approximation of the invariant measure to the stochastic equation (4.1).
Theorem 4.5.
Suppose that (4.2) and (4.3) hold. Let $s\geq 0$ and consider the particular time grid $t_{k}=s+\sum_{i=1}^{k}\frac{1}{i+1}$ for $k\in{\mathbb{N}}$. Then, there exists a constant $C\in{\mathbb{R}}_{+}$ such that we have for all $n\geq 1$,
$W_{1}^{|\cdot|}(\Theta_{s,t_{n}}^{\pi}(\nu),\mu_{\ast})\leq Cn^{-b_{\ast}\wedge 1/2}(1+\ln(1+n)^{\mathbf{1}_{b_{*}=1/2}})$ (4.7)
where $b_{\ast}$ is the constant in (4.5).
Proof.
Using Lemma 3.1 we get that $\Theta$ is $(1,b_{\ast},\frac{1}{2})$- coupled with $\theta.$ We already know that $\theta$ is $1$-bounded by Remark 4.2. It is also $(1,b_{\ast},\frac{1}{2})$-self coupled from (4.5) since we have $W_{1}(\theta_{s,t}(\mu),\theta_{s,t}(\nu))\leq W_{1}(\mu,\nu)(1-b_{\ast}(t-s)+\frac{b_{\ast}^{2}}{2}(t-s)^{2})$ and $W_{1}(\mu,\nu)\leq\Gamma_{1}(\mu,\nu)$. Then, by Theorem 2.12 and (4.4) we get (4.7). $\square$
Remark 4.6.
This analysis of the convergence of the Euler scheme has already been done in a more precise setting by Pagès and Panloup [PP23]. Here, we just want to illustrate how this enters in our abstract framework.
We give now a second example which is much more involved. Following Schuh [Sch24], we consider the space and time Langevin equation of McKean-Vlasov type
$dX_{t}$ $=$ $Y_{t}dt$ (4.8)
$dY_{t}$ $=$ $(ub^{E}(X_{t})+u\int_{{\mathbb{R}}^{d}}b^{I}(X_{t},z)\mu_{t}^{X}(dz)-\gamma Y_{t})dt+\sqrt{2\gamma u}dB_{t}$
where $\mu_{t}^{X}$ is the law of $X_{t},B$ is a $d$ dimensional Brownian motion, $b^{E}$ describes some external forces and $b^{I}$ represent some interacting forces. The parameters $u$ and $\gamma$ are positive real numbers associated to the physical model presented in [Sch24].
We assume that both $b^{E}$ and $b^{I}$ are Lipschitz continuous and moreover, we assume the following contraction property: There exists a positive definite matrix $K$ with smaller eigenvalue $\kappa>0$ and a Lipschitz continuous function $g:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ with Lipschitz constant $L_{g}$ such that $b^{E}(x)=-Kx+g(x)$. Moreover, the function $g$ verifies for some $R>0,$
$\left\langle x_{1}-x_{2},g(x_{1})-g(x_{2})\right\rangle\leq 0\quad\text{if}\quad\left|x_{1}-x_{2}\right|\geq R.$
The above assumption implies in particular that
$\left\langle x_{1}-x_{2},b^{E}(x_{1})-b^{E}(x_{2})\right\rangle\leq-\kappa\left|x_{1}-x_{2}\right|^{2}\quad\text{if}\quad\left|x_{1}-x_{2}\right|\geq R.$
Notice that the contraction property is not a classical one on $(X_{t},Y_{t})$, as the drift coefficient $b^{E}$ depends only the first component and it appears only on the second component. In this sense, we are dealing with a degenerate problem, and then the use of the distance $d_{f}$ from the previous example is no longer sufficient in order to solve that problem. Remarkably, Schuh constructs a new metric $d$ on ${\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ which is equivalent with the Euclidean distance (see [Sch24, Lemma 4.6]) and for which the following contraction result is obtained, see [Sch24, Theorem 12].
Theorem 4.7.
Suppose that the above assumptions hold with $2L_{g}^{2}u\gamma^{-2}<\kappa$. Let $\mu,\mu^{\prime}\in\mathcal{P}_{1}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$, $(X_{t},Y_{t})$ and $(X_{t}^{\prime},Y_{t}^{\prime})$ be two solutions starting respectively from $\mu$ and $\mu^{\prime}$. There exists a constant $b_{\ast}>0$ such that
$W_{1}(\mathcal{L}(X_{t},Y_{t}),\mathcal{L}(X_{t}^{\prime},Y_{t}^{\prime})))\leq e^{-b_{\ast}t}W_{1}(\mu,\mu^{\prime}).$ (4.9)
Moreover, there is a unique invariant measure for the corresponding semigroup $\theta$ and in particular, $\theta$ is $1$-bounded.
The last statement is a straightforward consequence of Proposition 2.11 and Remark 4.2.
The construction of the metric $d$ is rather heavy, but, as in the previous simpler example, we are not concerned with the specific form of the metric. Using Schuh’s contraction results for the semigroup is enough to directly obtain the needed conditions for the Euler scheme.
Remark 4.8.
Since $d$ is equivalent with the Euclidean distance, the above inequality also implies, for some $M>1$,
$W_{1}^{|\cdot|}(\mathcal{L}(X_{t},Y_{t}),\mathcal{L}(X_{t}^{\prime},Y_{t}^{\prime}))\leq Me^{-b_{*}t}W_{1}^{|\cdot|}(\mu,\mu^{\prime}),$
where $W_{1}^{|\cdot|}$ is the classical $1$-Wasserstein distance associated to the $1$-norm on ${\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$. However, this is no longer a strict contraction for $|\cdot|$ since $M>1$.
We define $\theta_{s,t}(\mu)$ as the law of the solution of (4.8) which starts from $\mu$ at time $s$. Let $\Theta_{s,t}(\mu$) be the law of $(\overline{X}_{s,t},\overline{Y}_{s,t})$, the one step Euler scheme defined by
$\overline{X}_{s,t}$ $=$ $\overline{Y}(t-s)$
$\overline{Y}_{s,t}$ $=$ $\left(ub^{E}(\overline{X})+u\int_{{\mathbb{R}}^{d}}b^{I}(\overline{X},z)\overline{\mu}^{\overline{X}}(dz)-\gamma\overline{Y}\right)(t-s)+\sqrt{2\gamma u}(B_{t}-B_{s})$
where $(\overline{X},\overline{Y})$ has law $\mu$ and $\mu^{\overline{X}}(dz)$ is the law of $\overline{X}$.
Theorem 4.9.
Let the assumptions of Theorem 4.7 be in force. Let $s\geq 0$ and $t_{n}=s+\sum_{i=1}^{n}\frac{1}{i+1}$ for $n\in{\mathbb{N}}$. Then, there exists $C\in{\mathbb{R}}_{+}$, such that for all $n\geq 1$
$W_{1}^{|\cdot|}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq Cn^{-\frac{1}{2}\wedge b_{\ast}}(1+\ln(n+1)^{1_{\{b_{\ast}=\frac{1}{2}\}}})$ (4.10)
Proof.
Let $(X_{t},Y_{t})$ be the solution starting from $(\overline{X},\overline{Y})$ such that $\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)<\infty$. We claim that there exists a constant $C\in{\mathbb{R}}_{+}$ such that for $s\leq t\leq s+1$,
$\mathbb{E}\left|X_{s,t}-\overline{X}_{s,t}\right|+\mathbb{E}\left|Y_{s,t}-\overline{Y}_{s,t}\right|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{3/2}.$ (4.11)
We first prove that
$\mathbb{E}\left|X_{s,t}\right|+\mathbb{E}\left|Y_{s,t}\right|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right).$ (4.12)
To get this, we use the triangle inequality in (4.8), observe that $\sqrt{2\gamma u}\mathbb{E}[|B_{t}-B_{s}|]$ is bounded for $t\leq s+1$, and conclude with Gronwall lemma.
Then, we use the Lipschitz property (and thus the sublinear growth) of $b^{E}$ and $b^{I}$. Using (4.12), this gives for $t\leq s+1$,
$\mathbb{E}|X_{s,t}-\overline{X}|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s),$
$\mathbb{E}|Y_{s,t}-\overline{Y}|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)+C(t-s)^{1/2}.$
With these estimates, we can finally compare the solution $X$ with the Euler scheme $\bar{X}$ and get
$\mathbb{E}(\left|X_{s,t}-\overline{X}_{s,t}\right|)\leq\int_{s}^{t}\mathbb{E}(\left|Y_{s,r}-\overline{Y}\right|)dr\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{2}+C(t-s)^{3/2},$
and, using the Lipschitz property of $b^{E}$ and $b^{I}$,
$\mathbb{E}(\left|Y_{s,t}-\overline{Y}_{s,t}\right|)\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{2}+C(t-s)^{3/2}.$
This gives directly (4.11) since $t-s\leq 1$.
From the equivalence between the distance $d$ and the Euclidean norm, we also have
$\mathbb{E}(d(X_{s,t},\overline{X}_{s,t}))+\mathbb{E}(d(Y_{s,t},\overline{Y}_{s,t}))\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{3/2},$
and therefore for any $\mu\in\mathcal{P}_{1}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$,
$W_{1}(\Theta_{s,t}(\mu),\theta_{s,t}(\mu))\leq C(1+\|\mu\|_{1})(t-s)^{3/2}.$
Moreover, using the triangle inequality and (4.9), we get
$W_{1}(\Theta_{s,t}(\nu),\theta_{s,t}(\mu))\leq C(1+\|\nu\|_{1})(t-s)^{3/2}+e^{-b_{*}t}W_{\rho,1}(\mu,\nu).$
So, $\theta$ and $\Theta$ are $(1,b_{*},1/2)$-coupled and (4.10) is then a consequence of Theorem 2.12. ∎
5. Examples with global contraction conditions
In this section, we propose to investigate different examples of stochastic systems that fall into the framework presented in Section 2. In contrast with Section 4 where the coupling property between the approximation scheme and the flow is deduced from the self-coupling property of the flow, we instead present in these section cases where the coupling property is deduced from the self-coupling property of the approximation scheme. When applying it we will need to prove the main coupling properties as described in Definition 2.3, as well as the $p$-boundedness stated in (2.4). In most examples, the main argument is based on an application of Theorem 3.7. Therefore the proofs consist in the verification of (3.4) and (2.22). The first example on reflected SDEs considers a rather simple example with Markovian dynamics without law dependence on the coefficients. Then, we consider a neuronal model, which is an example of McKean-Vlasov SDE. To show that the general framework in Section 2 can be applied in Wasserstein spaces with different value of $p$, we will obtain conditions which imply the convergence of the approximation schemes to the invariant measure in $W_{2}$ and $W_{1}$. This will lead to different contraction hypothesis. In these two examples, we can apply directly the method presented in Subsections 2.3 and 3.2. Namely, we construct a coupling between the continuous stochastic process and its approximation scheme. In contrast, for the Boltzmann equation approximation in $W_{2}$, we are not able to construct a coupling between the Boltzmann process and its one-step approximation scheme. To get around this difficulty, we will couple an approximation of the Boltzmann process (with say parameter $m$) with the one-step approximation scheme and then take limits as $m\to\infty$ in all the estimates. For the Boltzmann equation in $W_{1}$, we can skip this step relying on the sewing lemma. Note that in all the examples, we have not strived to find the best conditions as our goal is to show in a compact form that the main framework applies in each situation. Considering weaker conditions or more general equations in each example may be possible, but outside the scope of the paper. For readability, we have selected some proofs to keep them in the main text while the other ones are available in Appendix B.
Throughout this section, we assume that the time grid is taken so that $t_{n}=t_{0}+\sum_{i=1}^{n}\frac{1}{i+1}.$ The flows are always time homogeneous and continuous.
5.1. A stochastic equation with reflection in $W_{2}$
We consider the following one dimensional stochastic equation with reflection in $\mathbb{R}_{+}$ for the initial condition $X^{1}\geq 0$, $t-s\leq 1$ on a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ supporting a Wiener process $W$:
$\mathcal{X}^{1}_{s,t}(X^{1})=$ $X^{1}+\int_{s}^{t}b(\mathcal{X}^{1}_{s,u}(X^{1}))du+\int_{s}^{t}\sigma(\mathcal{X}^{1}_{s,u}(X^{1}))dW_{u}+L^{1}_{s,t}\geq 0,$ (5.1)
$L^{1}_{s,t}=$ $\int_{0}^{t}1_{\mathcal{X}^{1}_{s,u}(X^{1})=0}dL^{1}_{s,u}.$ (5.2)
Here, a solution means a couple $(\mathcal{X}^{1}_{s,\cdot},L^{1}_{s,\cdot})$ which satisfies the above conditions, where $L^{1}_{s,\cdot}$ is a nondecreasing process which only increases when $\mathcal{X}^{1}_{s,u}(X^{1})=0$ with initial condition $L^{1}_{s,s}=0$. We assume that $b,\sigma:\mathbb{R}\to\mathbb{R}$ are Lipschitz functions. In fact, we assume that there exists constants $C_{b},C_{\sigma}>0$ such that
$|b(y)-b(x)|^{2}\leq C_{b}|y-x|^{2},\quad|\sigma(y)-\sigma(x)|^{2}\leq C_{\sigma}|y-x|^{2}.$ (5.3)
Therefore, there exists a unique solution to the above equation, see for example [LS84, Theorem 3.1], and this solution has finite moments. In particular, we have ${\mathbb{E}}[\mathcal{X}^{1}_{s,t}(X^{1})^{2}]\leq C(1+{\mathbb{E}}[(X^{1})^{2}])$. We also define for $X^{2}\geq 0$, the one step approximation process
$\mathcal{X}_{s,t}^{2}(X^{2})=X^{2}+b(X^{2})(t-s)+\sigma(X^{2})(W_{t}-W_{s})+L_{s,t}^{2}\geq 0,$ (5.4)
with $L^{2}$ a nondecreasing process such that $L_{s,t}^{2}=\int_{s}^{t}1_{\mathcal{X}_{s,u}^{2}(X^{2})=0}dL_{s,u}^{2}$ with initial condition $L^{2}_{s,s}=0$. We similarly have ${\mathbb{E}}[\mathcal{X}^{2}_{s,t}(X^{2})^{2}]\leq C(1+{\mathbb{E}}[(X^{2})^{2}])$.
Furthermore, we assume the following contraction property: for some positive constant $\bar{b}>C_{\sigma}$
$(x-y)(b(x)-b(y))\leq-\bar{b}\left|x-y\right|^{2}.$ (5.5)
We intend to prove that the probabilistic representations implied by $\mathcal{X}^{i}$, $i=1,2$ satisfy the conditions in Theorem 2.15. In fact, we define for $\mu=\mathcal{L}(X)\in\mathcal{P}_{2}(\mathbb{R})$
$\theta_{s,t}(\mu)=\mathcal{L}(\mathcal{X}^{1}_{s,t}(X)),\quad\Theta_{s,t}(\mu)=\mathcal{L}(\mathcal{X}_{s,t}^{2}(X)).$
We fix now the time grid $\pi=\{t_{n},n\in{\mathbb{N}}\}$, with $t_{n}=s+\sum_{i=1}^{n}\frac{1}{1+i}$ and define $\Theta_{s,t_{n}}^{\pi}(\mu)$ by (2.3). We also define the Euler scheme $X_{0}^{\pi}=X$ and $X_{k+1}^{\pi}=\mathcal{X}_{t_{k},t_{k+1}}^{2}(X_{k}).$ Then, if the law of $X_{0}$ is $\mu,$ then we have $\mathcal{L}(X_{n}^{\pi})=\Theta_{s,t_{n}}^{\pi}(\mu).$
Theorem 5.1.
Suppose that (5.3) and (5.5) hold with $2\bar{b}>C_{\sigma}$. Let $b_{\ast}\in(0,2\bar{b}-C_{\sigma})$. Then, $\theta$ and $\Theta$ satisfy the $2$-Foster-Lyapunov condition, are $(2,b_{\ast},1)$-coupled and $\theta$ is self-coupled with the same parameters.
Furthermore, the Markov process $(\mathcal{X}_{s,t}^{1}(X))_{t\geq s}$ admits a unique invariant measure $\nu\in\mathcal{P}_{2}(\mathbb{R}_{+})$, and there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}_{+})$, $n\geq 1$ and $t>s$,
$W_{2}^{2}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq$ $C(\left\|\mu\right\|_{2}^{2}+1)\frac{1}{n^{b_{*}\wedge 1}}(1+\ln(n+1)^{1_{b_{*}=1}}),$
$W_{2}^{2}(\theta_{s,t}(\mu),\nu)\leq$ $C(\left\|\mu\right\|_{2}^{2}+1)\times(1+(t-s)^{1_{b_{*}=1}})e^{-(b_{*}\wedge 1)(t-s)}.$
The proofs are based on upper bounds for the $L^{2}$ norms of $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{2}(X^{1})$ as well as $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{1}(X^{2})$. These are performed using Ito’s formula and then the hypotheses give that the joint probabilistic representation is satisfied as stated in Definition 2.14. We refer to Appendix B for the details.
5.2. A neuronal model
In this section, we illustrate how our general framework fits an example with dynamics of McKean-Vlasov type. The following neuronal model has been developed and studied in [DMGLP15],[FL16], and [CTV20]. Let $X$ be the unique solution of the following mean field type stochastic equation
$X_{t}=X_{0}+\int_{0}^{t}b(X_{s})ds+J\int_{0}^{t}{\mathbb{E}}[f(X_{s})]ds-\int_{0}^{t}\int_{\mathbb{R}_{+}}X_{u-}1_{z\leq f(X_{u}-)\wedge M}dN(u,z),$ (5.6)
where $N$ is a random Poisson measure [^3] We have preferred the notation $dN(u,z)$ instead $N(du,dz)$ for random Poisson measures in order to get shorter formulas, but we admit that both notations are equivalent. with compensator measure given by the Lebesgue measure on $\mathbb{R}_{+}\times\mathbb{R}_{+}$. Furthermore, $M>0$ and $b,f:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ satisfy:
$\forall x,y\geq 0,|b(x)-b(y)|\leq C_{b}|x-y|,\ |f(x)-f(y)|\leq C_{f}|x-y|\text{ and }f(0)=0.$ (5.7)
In the case that $\|f\|_{\infty}\leq M$, we recover the model in [CTV20].
Let us define $f_{M}(x)=f(x)\wedge M$. Then, for $X^{i}\geq 0$, we define
$\mathcal{X}^{i}_{s,t}(X^{i})=$ $X^{i}+\int_{s}^{t}b_{i}(u)du+J\int_{s}^{t}{\mathbb{E}}[f_{i}(u)]du$
$-\int_{s}^{t}\int_{\mathbb{R}_{+}}\mathcal{X}^{i}_{s,u-}(X^{i})1_{z\leq f_{M}(\mathcal{X}^{i}_{s,u-}(X^{i}))}dN(u,z),$ (5.8)
and
$(b_{i}(u),f_{i}(u))=\begin{cases}(b(\mathcal{X}^{1}_{s,u}(X^{1})),f(\mathcal{X}^{1}_{s,u}(X^{1})));&\text{ if }i=1,\\ (b(X^{2}),f(X^{2}));&\text{ if }i=2.\end{cases}$
We assume moreover the following contraction condition for $x,y\in\mathbb{R}_{+}$
$2(y-x)(b(y)-b(x))\leq-\bar{b}(y-x)^{2}+y^{2}f_{M}(y)+x^{2}f_{M}(x)-2xyf_{M}(y)\vee f_{M}(x).$ (5.9)
This condition combines drift and jump effects in the model to create the contraction.
Remark 5.2.
Let us note that $\mathcal{X}^{2}_{s,t}$ stands for the approximation of the continuous process $\mathcal{X}^{1}_{s,t}$ but is different from the Euler scheme. Indeed, the coefficient of the driving Poisson measure depends on $\mathcal{X}^{2}_{s,u-}(X^{2})$, not on $X_{2}$. This plays a crucial role in the analysis below, in particular for the term (5.14). Nonetheless, the process $\mathcal{X}^{2}_{s,t}(X^{2})$ can be sampled exactly, as a piecewise deterministic Markov process.
From the Lipschitz condition 5.7, there exists a constant $K>0$ (which depends on $b(0)$) such that $|b(x)|\leq K(1+|x|)$. We define the following maps for $\mu=\mathcal{L}(X)\in\mathcal{P}_{2}(\mathbb{R}_{+})$
$\theta_{s,t}(\mu)=\mathcal{L}(\mathcal{X}^{1}_{s,t}(X)),\quad\Theta_{s,t}(\mu)=\mathcal{L}(\mathcal{X}_{s,t}^{2}(X)).$
Theorem 5.3.
Assume that (5.7) and (5.9) hold with $\bar{b}-2C_{f}J>0$. Then, $\theta$ and $\Theta$ satisfy the $2$-Foster-Lyapunov condition, are $(2,b_{\ast},1)$ coupled for every $b_{\ast}<\bar{b}-2C_{f}J$, and $\theta$ is self-coupled with the same parameters.
Furthermore, the Markov process $(\mathcal{X}_{s,t}^{1}(X))_{t\geq s}$ admits a unique invariant measure $\nu\in{\mathcal{P}}_{2}({\mathbb{R}_{+}})$, and for any $\beta\in(0,1]\cap(0,\bar{b}-2C_{f}J)$ there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in{\mathcal{P}}_{2}({\mathbb{R}_{+}})$, $n\geq 1$ and $t>s$,
$W_{2}^{2}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq$ $C(\left\|\mu\right\|_{2}^{2}+1)\frac{1}{n^{\beta}},$
$W_{2}^{2}(\theta_{s,t}(\mu),\nu)\leq$ $C(\left\|\mu\right\|_{2}^{2}+1)e^{-\beta(t-s)},$
where $t_{k}=s+\sum_{i=1}^{k}\frac{1}{1+i}$ and $\pi=\{t_{0}<\dots<t_{n}\}$.
Proof.
The proof consists in proving the conditions (3.4) and (2.22) and then apply Theorem 3.7.
Given the assumptions, we observe that the first two moments of $\mathcal{X}^{i}(X^{i})$, $i=1,2$ are bounded. In fact, for $t-s\leq 1$ applying Itô’s formula for $\mathcal{X}^{i}_{s,t}(X^{i})^{2}$, one obtains
$\mathcal{X}^{i}_{s,t}(X^{i})^{2}-(X^{i})^{2}=$ $2\int_{s}^{t}(\mathcal{X}^{i}_{s,u}(X^{i})(b_{i}(u)+J{\mathbb{E}}[f_{i}(u)]))du$
$-\int_{s}^{t}\int_{\mathbb{R}_{+}}\mathcal{X}^{i}_{s,u-}(X^{i})^{2}1_{z\leq f_{M}(\mathcal{X}^{i}_{s,u-}(X^{i}))}dN(u,z).$
Taking expectations in the above expression and evaluating the space integral, we obtain
${\mathbb{E}}[\mathcal{X}^{i}_{s,t}(X^{i})^{2}]-{\mathbb{E}}[(X^{i})^{2}]$
$=2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{i}_{s,u}(X^{i})(b_{i}(u)+J{\mathbb{E}}(f_{i}(u)))]du-\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{i}_{s,u}(X^{i})^{2}f_{M}(\mathcal{X}^{1}_{s,u}(X^{i}))]du$ (5.10)
$\leq 2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{i}_{s,u}(X^{i})(b_{i}(u)+J{\mathbb{E}}(f_{i}(u)))]du.$
Applying the inequality $2xy\leq x^{2}+y^{2}$, the Lipschitz properties of $b$ and $f$ one obtains from Gronwall’s lemma that
${\mathbb{E}}[\mathcal{X}^{i}_{s,t}(X^{i})^{2}]\leq C(1+{\mathbb{E}}[(X^{i})^{2}]).$ (5.11)
Here, $C$ depends on $C_{b}$, $C_{f}$ and $b(0)\geq 0$. Furthermore, using the equality (5.10) with the inequality (5.11), we get
$\left|{\mathbb{E}}[\mathcal{X}^{i}_{s,t}(X^{i})^{2}]-{\mathbb{E}}[(X^{i})^{2}]\right|\leq C(1+{\mathbb{E}}[(X^{i})^{2}])(t-s).$ (5.12)
The constant $C$ depends on $b(0)$, $C_{b}$, and $JC_{f}$ and $M$.
Using the contraction hypothesis (5.9) for $x=0$, $f_{M}(0)=f(0)=0$ and the inequality $2yb(0)\leq\delta y^{2}+\delta^{-1}b(0)^{2}$, we obtain the following inequality for any $\delta>0$
$2yb(y)-y^{2}f_{M}(y)\leq-(\bar{b}-\delta)y^{2}+\delta^{-1}b(0)^{2}.$
Therefore using the Lipschitz property of the positive function $f$ and Cauchy-Schwarz inequality, we obtain
$2{\mathbb{E}}[\mathcal{X}^{1}_{s,u}(X^{1})b(\mathcal{X}^{1}_{s,u}(X^{1}))]-{\mathbb{E}}[\mathcal{X}^{1}_{s,u}(X^{1})^{2}f_{M}(\mathcal{X}^{1}_{s,u}(X^{1}))]+2J{\mathbb{E}}[\mathcal{X}^{1}_{s,u}(X^{1})f(\mathcal{X}^{1}_{s,u}(X^{1}))]$
$\leq$ $-(\bar{b}-\delta){\mathbb{E}}[\mathcal{X}^{1}_{s,u}(X^{1})^{2}]+\delta^{-1}b(0)^{2}+2JC_{f}\mathbb{E}[\mathcal{X}^{1}_{s,u}(X^{1})^{2}].$
Therefore, applying the above inequality in (5.10) for $i=1$, we obtain
${\mathbb{E}}[\mathcal{X}^{1}_{s,t}(X^{1})^{2}]-{\mathbb{E}}[(X^{1})^{2}]\leq$ $\int_{s}^{t}-(\bar{b}-\delta-2JC_{f}){\mathbb{E}}[\mathcal{X}^{1}_{s,u}(X^{1})^{2}]du+\delta^{-1}b(0)^{2}(t-s).$
Finally applying again (5.12), one obtains the Foster Lyapunov criterion (3.4) for $p=2$ with ${\varepsilon}=1$, $\widetilde{C}=\delta^{-1}b(0)^{2}$, $\hat{C}=C(\bar{b}-\delta-2JC_{f})$ and $\hat{b}=\bar{b}-\delta-2JC_{f}$. Here $C$ depends on $b(0)$, $C_{b}$, and $JC_{f}$ and $M$.
The proof for $\mathcal{X}^{2}$ is similar but requires a small change as follows. Decompose the right hand side of (5.10) for $i=2$ as follows:
$2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{2}_{s,u}(X^{2})(b(X^{2})-b(\mathcal{X}^{2}_{s,u}(X^{2}))]+J{\mathbb{E}}[f(X^{2})-f(\mathcal{X}^{2}_{s,u}(X^{2}))]du$
$+2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{2}_{s,u}(X^{2})(b(\mathcal{X}^{2}_{s,u}(X^{2}))+J{\mathbb{E}}[f(\mathcal{X}^{2}_{s,u}(X^{2}))]du$
$-\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{2}_{s,u}(X^{2})^{2}f_{M}(\mathcal{X}^{2}_{s,u}(X^{2}))]du$
The last two integrals are treated as in the case for $\mathcal{X}^{1}_{s,u}(X^{1})$, while the first integral is estimated using Cauchy-Schwarz inequality, the Lipschitz property (5.7) and the moment estimate (5.12) together with [^4] We tame the jump intensity in (5.6) by $M$ precisely to get this bound. Otherwise the (negative) contribution of jumps may not be sufficiently integrable.
${\mathbb{E}}|\mathcal{X}^{2}_{s,u}(X^{2})-X^{2}|^{2}\leq C(1+{\mathbb{E}}(X^{2})^{2})(u-s).$
Here $C$ depends on $b(0)$, $C_{b}$, $JC_{f}$ and $M$. This gives (3.4) with $\varepsilon=1/2$ (due to the Cauchy-Schwarz inequality that was not used for $\mathcal{X}^{1}$ in the term $J{\mathbb{E}}[f(X^{2})-f(\mathcal{X}^{2}_{s,u}(X^{2}))]$) and a slightly more involved $\hat{C}$ due to the extra estimates required. The constants $\widetilde{C}$ and $\hat{b}$ remained unchanged.
Next, we consider the difference $\mathcal{Y}:=\mathcal{X}^{1}(X^{1})-\mathcal{X}^{2}(X^{2})$. Define
$\Delta b(u):=$ $b(\mathcal{X}^{1}_{s,u}(X^{1}))-b(X^{2})$
$\Delta{\mathbb{E}}[f](u):=$ ${\mathbb{E}}[f(\mathcal{X}^{1}_{s,u}(X^{1}))]-{\mathbb{E}}[f(X^{2})].$
Again, compensating the jump terms and applying Itô’s formula, we have
${\mathbb{E}}[\mathcal{Y}_{s,t}^{2}]={\mathbb{E}}(X^{1}-X^{2})^{2}+2\int_{s}^{t}{\mathbb{E}}[\mathcal{Y}_{s,u}\left(\Delta b(u)+J\Delta{\mathbb{E}}[f](u)\right)]du$
$-2\int_{s}^{t}{\mathbb{E}}\left[\mathcal{Y}_{s,u}(\mathcal{X}^{1}_{s,u}(X^{1})f_{M}(\mathcal{X}^{1}_{s,u}(X^{1}))-\mathcal{X}^{2}_{s,u}(X^{2})f_{M}(\mathcal{X}^{2}_{s,u}(X^{2})))\right]du$ (5.13)
$+\int_{s}^{t}\int_{\mathbb{R}_{+}}{\mathbb{E}}\left[\left(\mathcal{X}^{1}_{s,u-}(X^{1})1_{z\leq f_{M}(\mathcal{X}^{1}_{s,u-}(X^{1}))}-\mathcal{X}^{2}_{s,u-}(X_{2})1_{z\leq f_{M}(\mathcal{X}^{2}_{s,u-}(X^{2}))}\right)^{2}\right]dzdu.$ (5.14)
From the above formula, we can use the boundedness of moments of $\mathcal{X}^{i}(X^{i})$ to obtain that
$\left|{\mathbb{E}}[\mathcal{Y}_{s,t}^{2}]-{\mathbb{E}}[(X^{1}-X^{2})^{2}]\right|\leq C(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(t-s).$ (5.15)
With the above preparation, we proceed with the proof of (2.22). We start with an estimate for the last integral in (5.14). Evaluating the space integral, this term can be rewritten as
$\int_{s}^{t}{\mathbb{E}}\big{[}(\mathcal{X}^{1}_{s,u}(X^{1}))^{2}f_{M}(\mathcal{X}^{1}_{s,u}(X^{1}))-2\mathcal{X}^{1}_{s,u}(X^{1})\mathcal{X}^{2}_{s,u}(X_{2})f_{M}(\mathcal{X}^{1}_{s,u}(X^{1}))\wedge f_{M}(\mathcal{X}^{2}_{s,u}(X^{2}))$
$+(\mathcal{X}^{2}_{s,u}(X_{2}))^{2}f_{M}(\mathcal{X}^{2}_{s,u}(X^{2}))\big{]}du.$
Now consider the above expression together with (5.13) and replace $(\mathcal{X}^{1}_{s,u}(X^{1}),\mathcal{X}^{2}_{s,u}(X^{2}))$ with $(y,x)$, then one obtains the following expression:
$-2(y-x)(yf_{M}(y)-xf_{M}(x))+y^{2}f_{M}(y)-2xyf_{M}(y)\wedge f_{M}(x)+x^{2}f_{M}(x)$
$=-y^{2}f_{M}(y)-x^{2}f_{M}(x)+2xyf_{M}(x)\vee f_{M}(y)\leq-\bar{b}(y-x)^{2}-2(y-x)(b(y)-b(x)),$
by using the contraction inequality (5.9). Therefore, putting the above estimates together we have
${\mathbb{E}}[\mathcal{Y}_{s,t}^{2}]-{\mathbb{E}}[(X^{1}-X^{2})^{2}]\leq$ $\int_{s}^{t}-\bar{b}{\mathbb{E}}[\mathcal{Y}_{s,u}^{2}]+2{\mathbb{E}}[\mathcal{Y}_{s,u}(b(\mathcal{X}^{2}_{s,u}(X^{2}))-b(X^{2}))]$
$+2J{\mathbb{E}}[\mathcal{Y}_{s,u}({\mathbb{E}}f(\mathcal{X}^{1}_{s,u}(X^{1}))-{\mathbb{E}}f(\mathcal{X}^{2}_{s,u}(X^{2})))]$
$+2J{\mathbb{E}}[\mathcal{Y}_{s,u}({\mathbb{E}}f(\mathcal{X}^{2}_{s,u}(X^{2}))-{\mathbb{E}}f(X^{2}))]du$
$\leq$ $\int_{s}^{t}-(\bar{b}-\delta-2JC_{f}){\mathbb{E}}\mathcal{Y}_{s,u}^{2}du+C\delta^{-1}{\mathbb{E}}(1+|X^{2}|^{2})(t-s)^{2}.$
Again we have used the inequality $2xy\leq\delta x^{2}+\delta^{-1}y^{2}$ the Lipschitz property of $b$ and $f$ and the moment estimate (5.12). Therefore, using (5.15), we get
${\mathbb{E}}[\mathcal{Y}_{s,t}^{2}]\leq\left(1-(\bar{b}-\delta-2JC_{f})\right){\mathbb{E}}[(X^{1}-X^{2})^{2}]+C(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(t-s)^{2},$ (5.16)
so that (2.22) is satisfied for any $b_{*}\in(0,\bar{b}-2JC_{f})$ with $\varepsilon=1$ and $p=2$.
The proof for the self-coupling property for $\theta$ is similar but simpler. ∎
5.2.1. Neural model in $W_{1}$.
We now consider the same neuronal model (5.6) with the assumption (5.7). Instead of (5.9), we make the following contraction assumption: there exists $\bar{b}>0$ such that for any $x,y\geq 0$,
$\mathrm{sgn}(y-x)(b(y)-b(x))-|y-x|(f_{M}(y)\vee f_{M}(x))$
$+x(f_{M}(y)-f_{M}(x))^{+}+y(f_{M}(x)-f_{M}(y))^{+}\leq-\bar{b}\left|x-y\right|$ (5.17)
This allows to deal with only integrable distributions and to work with the $W_{1}$ distance. We also remark that in the particular case that $f_{M}$ is an non-decreasing function, we have
$-|y-x|(f_{M}(y)\vee f_{M}(x))+x(f_{M}(y)-f_{M}(x))^{+}+y(f_{M}(x)-f_{M}(y))^{+}\leq 0.$
This means that in some situations the drift does not need to have a contraction property.
We have under these assumptions the following result.
Theorem 5.4.
Assume that (5.7) and (5.17) hold with $\bar{b}-C_{f}J>0$. Then, $\theta$ and $\Theta$ satisfy the $1$-Foster-Lyapunov condition, are $(1,b_{*},1)$ coupled for every $b_{*}=\bar{b}-C_{f}J$, and $\theta$ is self-coupled with the same parameters.
Furthermore, the Markov process $(\mathcal{X}_{s,t}^{1}(X))_{t\geq s}$ admits a unique invariant measure $\nu\in{\mathcal{P}}_{1}({\mathbb{R}_{+}})$, and there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in{\mathcal{P}}_{1}({\mathbb{R}_{+}})$, $n\geq 1$ and $t>s$,
$W_{1}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq C(\left\|\mu\right\|_{1}+1)\frac{(1+\ln(n+1))^{\mathbf{1}_{b_{*}=1}}}{n^{b_{*}\wedge 1}},$
$W_{1}(\theta_{s,t}(\mu),\nu)\leq C(\left\|\mu\right\|_{1}+1)\times(1+(t-s)^{\mathbf{1}_{b_{*}=1}})e^{-(b_{*}\wedge 1)(t-s)},$
where $t_{k}=s+\sum_{i=1}^{k}\frac{1}{1+i}$ and $\pi=\{t_{0}<\dots<t_{n}\}$.
The proof of this result is detailed in the additional material in Appendix B. A slight difference with the $W_{2}$ case is that we are able to prove the contraction property for $b_{*}=\bar{b}-C_{f}J$ (instead of any $b_{*}\in(0,\bar{b}-C_{f}J)$), and therefore the special case $b_{*}=\varepsilon$ may occur.
5.3. Boltzmann type equations
In this section, we present two examples based on Boltzmann type equations. These are stochastic equations driven by a Poisson point measure whose compensator depends on the law of the solution of the equation. Such equations appear in the modelisation of interacting physical particles.
As in previous examples, we consider two different settings. In the first one, we work with Poisson point measures that corresponds to the “finite variation regime”. This is a simpler framework, which allows to consider a more complex dependence of the Poisson point measure on the position of the particle. This is the case of the Boltzmann equation with hard potentials for example (here we deal with a drastically simplified example). In order to be able to deal with the indicator function which appears in the stochastic integral we have to use $L^{1}$ estimates, and the appropriate distance here is $W_{1}.$
In the second setting, we are in the “martingale regime” and so we have to work with a compensated Poisson point measure. In this case, all the estimates have to be done in $L^{2}$ and we use the Wasserstein distance $W_{2}.$ This corresponds to Maxwellian molecules. Notice that in this case, no indicator function appears in the jump part of the equation, since we are not able to handle this type of discontinuity in the martingale regime. Furthermore the estimates in $W_{2}$ require different arguments as in the $W_{1}$ case because we cannot use Lemma 3.1 to get the coupling property.
5.3.1. The finite variation regime ($W_{1}$)
We consider the following $d$-dimensional stochastic equation
$X_{s,t}^{1}(X^{1})$ $=$ $X^{1}+\int_{s}^{t}b(X_{s,r}^{1}(X^{1}))dr$
$+\int_{s}^{t}\int_{E}\int_{\mathbb{R}_{+}}\int_{\mathbb{R}^{d}}c(v,z,X_{s,r-}^{1}(X^{1}))1_{\{u\leq\gamma(v,z,X_{s,r-}^{1}(X^{1}))\}}dN_{\mathcal{L}(X_{s,r}^{1})}(v,u,z,r)$
where $b:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$, $c:{\mathbb{R}}^{d}\times E\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and $dN_{\mathcal{L}(X_{s,r}^{1})}(v,u,z,r)$ is the Poisson point measure on $\mathbb{R}^{d}\times\mathbb{R}_{+}\times E\times\mathbb{R}_{+}$ with compensator $d\widehat{N}_{\mathcal{L}(X_{s,r}^{1})}(v,u,z,r)=\mathcal{L}(X_{s,r}^{1})(dv)du\nu(dz)dr.$ Here, $\nu$ is a measure on the measurable space $E$
We associate to the above equation, the one step Euler scheme
$X_{s,t}^{2}(X^{2})=X^{2}+\int_{s}^{t}b(X^{2})dr+\int_{s}^{t}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}_{+}}\int_{E}c(v,z,X^{2})1_{\{u\leq\gamma(v,z,X^{2}))\}}dN_{\mathcal{L}(X^{2})}(v,u,z,r)$ (5.19)
where $dN_{\mathcal{L}(X^{2})}(v,u,z,r)$ is the Poisson point measure with compensator $d\widehat{N}_{\mathcal{L}(X^{2})}(v,u,z,r)=\mathcal{L}(X^{2})(dv)du\nu(dz)dr.$
Moreover, we define the maps on the Wasserstein space for $\mathcal{L}(X)=\mu$
$\theta_{s,t}(\mu)=\mathcal{L}(X_{s,t}^{1}(X)),\quad\Theta_{s,t}(\mu)=\mathcal{L}(X_{s,t}^{2}(X)).$ (5.20)
We will work under the following hypotheses that guarantee that the equation (5.3.1) has a weak solution and that $\theta_{s,t}$ is well defined.
H1. For every $x,y\in\mathbb{R}^{d}$
$\left\langle x-y,b(x)-b(y)\right\rangle\leq-\overline{b}\left|x-y\right|^{2},\quad\left|b(x)-b(y)\right|\leq C_{b}\left|x-y\right|$ (5.21)
H2. For every $x,x^{\prime},v,v^{\prime}\in\mathbb{R}^{d}$ and $z\in E$ there exists a measurable function $\bar{c}:E\to\mathbb{R}_{+}$ such that $Q:=\int_{E}\overline{c}(z)\nu(dz)<\infty$ and
$\left|(c(v,z,x)-c(v^{\prime},z,x^{\prime}))\gamma(v,z,x)\right|+$ $\left|c(v^{\prime},z,x^{\prime})(\gamma(v,z,x)-\gamma(v^{\prime},z,x^{\prime}))\right|$
$\leq\overline{c}(z)(\left|v-v^{\prime}\right|+\left|x-x^{\prime}\right|)$ (5.22)
$\left|c\gamma(v,z,x)\right|\leq$ $\overline{c}(z)(1+\left|v\right|+\left|x\right|).$ (5.23)
Lemma 5.5.
Under Assumption H2, we have
$\int_{E}\int_{{\mathbb{R}}_{+}}\left|c(v,z,x)1_{u\leq\gamma(v,z,x)}-c(v^{\prime},z,x^{\prime})1_{u\leq\gamma(v^{\prime},z,x^{\prime})}\right|du\nu(dz)\leq Q(|v-v^{\prime}|+|x-x^{\prime}|).$
Proof.
The integral is upper bounded by
$\int_{E}\int_{{\mathbb{R}}_{+}}\left|c(v,z,x)-c(v^{\prime},z,x^{\prime})\right|1_{u\leq\gamma(v,z,x)}du\nu(dz)$
$+\int_{E}\int_{{\mathbb{R}}_{+}}|c(v^{\prime},z,x^{\prime})|\left|1_{u\leq\gamma(v,z,x)}-1_{u\leq\gamma(v^{\prime},z,x^{\prime})}\right|du\nu(dz)$
$=$ $\int_{E}\left|c(v,z,x)-c(v^{\prime},z,x^{\prime})\right|\gamma(v,z,x)+|c(v^{\prime},z,x^{\prime})|\left|\gamma(v,z,x)-\gamma(v^{\prime},z,x^{\prime})\right|\nu(dz)$
$\leq$ $Q(|v-v^{\prime}|+|x-x^{\prime}|).\qed$
Under hypotheses H1 and H2, we know by [AB23, Theorem 3.5] that there exists a unique time homogeneous continuous flow $\theta_{s,t}:\mathcal{P}_{1}({\mathbb{R}}^{d})\to\mathcal{P}_{1}({\mathbb{R}}^{d})$ such that
$\forall T>0,\exists C>0,\forall 0\leq s\leq t\leq T,\ \mathcal{D}_{1}(\theta_{s,t},\Theta_{s,t})\leq C(t-s)^{2},$ (5.24)
with $\mathcal{D}_{1}(\Theta,\overline{\Theta}):=\sup_{\mu\in\mathcal{P}_{1}({\mathbb{R}}^{d})}\frac{W_{1}(\Theta(\mu),\overline{\Theta}(\mu))}{1+\|\mu\|_{1}}$, and we also have
$W_{1}(\theta_{s,t}(\mu),\Theta_{t^{n}_{n},t^{n}_{n-1}}\circ\dots\circ\Theta_{t^{n}_{0},t^{n}_{1}}(\mu))\to_{n\to\infty}0$ (5.25)
for any sequence $s=t^{n}_{0}<t^{n}_{1}<\dots<t^{n}_{n}=t$ such that $\max_{1\leq i\leq n}t^{n}_{i}-t^{n}_{i-1}\to_{n\to\infty}0$. Besides, [AB23, Theorem 3.8] gives the existence of a solution to (5.3.1) such that $\mathcal{L}(X^{1}_{s,t}(X^{1}))=\theta_{s,t}(\mu^{1})$ when $\mathcal{L}(X^{1})=\mu^{1}\in\mathcal{P}_{1}({\mathbb{R}}^{d})$, and [AB23, Proposition 3.9] gives that any solution of (5.3.1) has the same marginal laws. We have the following result for the convergence of a scheme towards the invariant measure.
Theorem 5.6.
Assume that (5.21), (5.3.1) hold with $2Q<\overline{b}$. Then, $\theta$ and $\Theta$ satisfy the $1$-Foster-Lyapunov condition, are $(1,b_{*},1)$ coupled for $b_{*}=\bar{b}-2Q$, and $\theta$ is self-coupled with the same parameters.
Furthermore, the Markov process $({X}_{s,t}^{1}(X))_{t\geq s}$ admits a unique invariant measure $\nu\in{\mathcal{P}}_{1}({\mathbb{R}}^{d})$, and there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in{\mathcal{P}}_{1}({\mathbb{R}}^{d})$, $n\geq 1$ and $t>s$,
$W_{1}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq C(\left\|\mu\right\|_{1}+1)\frac{(1+\ln(n+1)^{\mathbf{1}_{b_{*}=1}})}{n^{b_{*}\wedge 1}},$
$W_{1}(\theta_{s,t}(\mu),\nu)\leq C(\left\|\mu\right\|_{1}+1)\times(1+(t-s)^{\mathbf{1}_{b_{*}=1}})e^{-b_{*}\wedge 1(t-s)},$
where $t_{k}=s+\sum_{i=1}^{k}\frac{1}{1+i}$ and $\pi=\{t_{0}<\dots<t_{n}\}$.
Proof.
The proof is an application of Theorem 3.7. From (5.24) and Lemma 3.1, it is sufficient to check that $\Theta$ is $(1,b_{*},1)$ self-coupled for $b_{*}=\overline{b}-2Q$, and satisfies the $1$-Foster-Lyapunov condition.
We consider a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ endowed with a Poisson point measure $dN(w,u,z,r)$ of intensity $dwdu\nu(dz)dr$ on $(0,1)\times\mathbb{R}_{+}\times E\times\mathbb{R}_{+}$. We denote by $(\mathcal{F}_{t})_{t}$ the filtration generated by this random measure. Moreover, we consider $X^{i}\in L^{1}(\Omega)$, $i=1,2$, which are $\mathcal{F}_{s}$ measurable and such that the distribution of $(X^{1},X^{2})$ is an $W_{1}$-optimal coupling for $\mu^{1}$ and $\mu^{2}.$ In particular $X^{i}\sim\mu^{i}$. We consider a measurable map $\tau:[0,1]\to{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ such that $\mathcal{L}(\tau(U))=\mathcal{L}(X^{1},X^{2})$ when $U$ is a $[0,1]$ uniform distributed random variable. We note $\tau^{i}$, $i=1,2$ its coordinates, and we define
$x_{s,t}^{i}=X^{i}+\int_{s}^{t}b(X^{i})dr+\int_{s}^{t}\int_{E}\int_{0}^{1}q(\tau^{i}(w),u,z,X^{i})dN(w,u,z,r)$ (5.26)
with
$q(v,u,z,x)=c(v,z,x)1_{\{u\leq\gamma(v,z,x)\}}.$
It is easy to check that $\mathcal{L}(x_{s,t}^{i})=\mathcal{L}(\Theta_{s,t}(\mu^{i}))$.
Next, we check the Foster-Lyapunov condition. Notice that, as an immediate consequence of our hypothesis (5.21) we have,
$\frac{1}{\left|x\right|}\left\langle x,b(x)\right\rangle=$ $\frac{1}{\left|x\right|}\left\langle x-0,b(x)-b(0)\right\rangle+\frac{1}{\left|x\right|}\left\langle x,b(0)\right\rangle\leq-\overline{b}\left|x\right|+\left|b(0)\right|$ (5.27)
$\left|b(x)\right|\leq$ $K(1+\left|x\right|).$ (5.28)
We then observe that
${\mathbb{E}}\left|x_{s,t}^{1}(X^{1})-X^{1}\right|$ $\leq$ $(t-s)\mathbb{E}[|b(X^{1})|]+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|c\gamma(\tau^{1}(w),z,X^{1})\right|dw\nu(dz)dr$ (5.29)
$\leq$ $(t-s)\left(K(1+\mathbb{E}[|X^{1}|])+Q(1+2\mathbb{E}[|X^{1}|])\right),$
using (5.28) and (5.23).
We now apply Itô’s formula with the function $F(x)=\left|x\right|$. In order to do this we consider $\psi_{\varepsilon}:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ defined by $\psi_{\varepsilon}(x)=\frac{x^{2}}{2\varepsilon}$ for $x\in[0,\varepsilon]$ and $\psi_{\varepsilon}(x)=x-\varepsilon/2$ for $x>\varepsilon$. We then define $F_{\varepsilon}(x)=\psi_{\varepsilon}(\left|x\right|)=\psi_{\varepsilon}(F(x))$: it is $C^{1}$ and we have $\lim_{\varepsilon\rightarrow 0}F_{\varepsilon}(x)=F(x)=\left|x\right|$ and $\lim_{\varepsilon\rightarrow 0}\partial_{i}F_{\varepsilon}(x)=\frac{x_{i}}{\left|x\right|}.$ We use now Itô’s formula (see e.g. [RW00, Theorem IV.18] for the finite variation case) and we get
${\mathbb{E}}[F_{\varepsilon}(x_{s,t}^{1})]$ $=$ ${\mathbb{E}}[F_{\varepsilon}(X^{1})]+{\mathbb{E}}\int_{s}^{t}\left\langle\nabla F_{\varepsilon}(x_{s,r}^{1}),b(X^{1})\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{\infty}\int_{0}^{1}(F_{\varepsilon}(x_{s,r}^{1}+q(\tau^{1}(w),u,z,X^{1}))-F_{\varepsilon}(x_{s,r}^{1}))dwdu\nu(dz)dr.$
Passing to the limit with $\varepsilon\rightarrow 0$ we get
${\mathbb{E}}[|x_{s,t}^{1}|]=$ ${\mathbb{E}}[|X^{1}|]+{\mathbb{E}}\int_{s}^{t}\frac{1}{|x_{s,r}^{1}|}\left\langle x_{s,r}^{1},b(X^{1})\right\rangle dr$ (5.30)
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{\infty}\int_{0}^{1}(\left|x_{s,r}^{1}+q(\tau^{1}(w),u,z,X^{1})\right|-|x_{s,r}^{1}|)dwdu\nu(dz)dr$
By using (5.27), we get
$\frac{1}{\left|x_{s,r}^{1}\right|}\left\langle x_{s,r}^{1},b(X^{1})\right\rangle$ $=$ $\frac{1}{\left|x_{s,r}^{1}\right|}\left\langle x_{s,r}^{1},b(x_{s,r}^{1})\right\rangle+\frac{1}{\left|x_{s,r}^{1}\right|}\left\langle x_{s,r}^{1},b(X^{1})-b(x_{s,r}^{1})\right\rangle$
$\leq$ $-\overline{b}\left|x_{s,r}^{1}\right|+\left|b(0)\right|+C_{b}\left|X^{1}-x_{s,r}^{1}\right|$
$\leq$ $-\overline{b}|X^{1}|+\left|b(0)\right|+(\overline{b}+C_{b})\left|X^{1}-x_{s,r}^{1}\right|.$
In particular, we get from (5.29)
${\mathbb{E}}\int_{s}^{t}\frac{1}{\left|x_{s,r}^{1}\right|}\left\langle x_{s,r}^{1},b(X^{1})\right\rangle dr\leq-\overline{b}(t-s){\mathbb{E}}[|X_{1}|]+\left|b(0)\right|(t-s)+C(1+{\mathbb{E}}\left|X^{1}\right|)(t-s)^{2}.$
We write now
${\mathbb{E}}\int_{s}^{t}\int_{0}^{1}\int_{E}\int_{\mathbb{R}_{+}}(\left|x_{s,r}^{1}+q(\tau^{1}(w),u,z,X^{1})\right|-\left|x_{s,r}^{1}\right|)dwdu\nu(dz)dr$
$=$ ${\mathbb{E}}\int_{s}^{t}\int_{0}^{1}\int_{E}\int_{\mathbb{R}_{+}}\left(\left|x_{s,r}^{1}+c(\tau^{1}(w),z,X^{1})\right|-|x_{s,r}^{1}|\right)1_{\{u\leq\gamma(\tau^{1}(w),z,X^{1})\}})dwdu\nu(dz)dr$
$\leq$ ${\mathbb{E}}\int_{s}^{t}\int_{0}^{1}\int_{E}\left|c\gamma(\tau^{1}(w),z,X^{1})\right|dw\nu(dz)dr$
$\leq$ $Q{\mathbb{E}}\int_{s}^{t}\int_{0}^{1}(1+|\tau^{1}(w)|+|X^{1}|)dw\nu(dz)dr=Q(t-s)(1+2{\mathbb{E}}[|X^{1}|]).$
Using the two last estimates in (5.30), we get
${\mathbb{E}}[\left|x_{s,t}^{1}\right|]\leq{\mathbb{E}}[\left|X^{1}\right|](1-(\overline{b}-2Q)(t-s))+(|b(0)|+Q)(t-s)+C(1+{\mathbb{E}}[\left|X^{1}\right|])(t-s)^{2},$
which proves the Foster-Lyapunov condition by Remark 3.5.
We now prove that $\Theta$ is $(1,b_{*},1)$-self-coupled. Let $y_{s,t}=x^{1}_{s,t}-x^{2}_{s,t}$. Using Itô’s formula (again, one has to take first a regularization and then to pass to the limit) we get
${\mathbb{E}}\left|y_{s,t}\right|$ $=$ ${\mathbb{E}}\left|X^{1}-X^{2}\right|+{\mathbb{E}}\int_{s}^{t}\left\langle\frac{y_{s,r}}{\left|y_{s,t}\right|},b(X_{1})-b(X_{2})\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{\mathbb{R}_{+}}\int_{0}^{1}\left|y_{s,r-}+\Delta q(w,u,z)\right|-\left|y_{s,r-}\right|dN(w,u,z,r)$
with $\Delta q(w,u,z)=q(\tau^{1}(w),u,z,X^{1})-q(\tau^{2}(w),u,z,X^{2})$. We will use this equality to do two different estimations.
First, we have
$\left|{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{\mathbb{R}_{+}}\int_{0}^{1}(\left|y_{s,r-}+\Delta q(r,u,w,z)\right|-\left|y_{s,r-}\right|)dN(w,u,z,r)\right|$
$\leq$ ${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{\mathbb{R}_{+}}\int_{0}^{1}\left|\Delta q(w,u,z)\right|dwdu\nu(dz)dr.$
By Lemma 5.5, we get
$\int_{E}\int_{\mathbb{R}_{+}}\left|\Delta q(w,u,z)\right|du\nu(dz)\leq Q\left(|\tau^{1}(w)-\tau^{2}(w)|+|X^{1}-X^{2}|\right),$
and thus
${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{\mathbb{R}_{+}}\int_{0}^{1}\left|\Delta q(r,u,w,z)\right|dwdu\nu(dz)dr\leq 2Q(t-s){\mathbb{E}}[|X^{1}-X^{2}|].$ (5.32)
Second, using that $b$ is Lipschitz, we obtain
$\left|{\mathbb{E}}\left|y_{s,t}\right|-{\mathbb{E}}\left|X^{1}-X^{2}\right|\right|\leq C(1+{\mathbb{E}}\left|X^{1}\right|+{\mathbb{E}}\left|X^{2}\right|)(t-s).$
In order to obtain (2.22), we perform a different estimation in the first integral term in (5.3.1). Notice that, by (5.21) and then (5.29), we have
${\mathbb{E}}\left\langle\frac{y_{s,r}}{\left|y_{s,t}\right|},b(X_{1})-b(X_{2})\right\rangle$ $\leq$ $-\overline{b}{\mathbb{E}}\left|y_{s,r}\right|+{\mathbb{E}}\left\langle\frac{y_{s,r}}{\left|y_{s,t}\right|},b(X^{1})-b(x_{s,r}^{1})\right\rangle+{\mathbb{E}}\left\langle\frac{y_{s,r}}{\left|y_{s,t}\right|},b(x_{s,r}^{2})-b(X^{2})\right\rangle$
$\leq$ $-\overline{b}{\mathbb{E}}\left|y_{s,r}\right|+C_{b}({\mathbb{E}}\left|x_{s,r}^{1}-X^{1}\right|+{\mathbb{E}}\left|x_{s,r}^{2}-X^{2}\right|)$
$\leq$ $-\overline{b}{\mathbb{E}}\left|y_{s,r}\right|+C(1+{\mathbb{E}}\left|X^{1}\right|+{\mathbb{E}}\left|X^{2}\right|)(r-s).$
We put the above estimate together with (5.32) in (5.3.1) and get
${\mathbb{E}}\left|y_{s,t}\right|$ $\leq{\mathbb{E}}\left|X^{1}-X^{2}\right|(1+2Q(t-s))-\overline{b}\int_{s}^{t}{\mathbb{E}}\left|y_{s,r}\right|dr+C(1+{\mathbb{E}}\left|X^{1}\right|+{\mathbb{E}}\left|X^{2}\right|)(t-s)^{2}$
$\leq$ $\left({\mathbb{E}}\left|X^{1}-X^{2}\right|(1+(2Q-\bar{b})(t-s))+C(1+{\mathbb{E}}\left|X^{1}\right|+{\mathbb{E}}\left|X^{2}\right|)(t-s)^{2}\right).$
This proves that $\Theta$ is $(1,b_{*},1)$-self-coupled. We conclude by using Lemma 3.1 and (3.1), the latter being given by (5.24). ∎
5.3.2. The martingale regime
We now consider the following $d$-dimensional stochastic equation for $X^{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$
$X_{s,t}^{1}(X^{1})=X^{1}+\int_{s}^{t}b(X_{s,r}^{1}(X^{1}))dr+\int_{s}^{t}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,X_{s,r-}^{1}(X^{1}))d\widetilde{N}_{\mathcal{L}(X_{s,r}^{1})}(v,z,r)$ (5.33)
where $b:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$, $c:{\mathbb{R}}^{d}\times E\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and $dN_{\mathcal{L}(X_{s,r}^{1})}(v,z,r)$ is a Poisson point measure on $\mathbb{R}^{d}\times E\times\mathbb{R}_{+}$ independent of $X_{1}$ with compensator $d\widehat{N}_{\mathcal{L}(X_{s,r}^{1})}(v,z,r)=\mathcal{L}(X_{s,r}^{1})(dv)\nu(dz)dr$ and $\widetilde{N}_{\mathcal{L}(X_{s,r}^{1})}=N_{\mathcal{L}(X_{s,r}^{1})}-\widehat{N}_{\mathcal{L}(X_{s,r}^{1})}.$ Here again, $\mathcal{L}(X_{s,r}^{1})$ denotes the law of the random variable $X_{s,r}^{1}$ and $E$ is a measure space with measure $\nu$.
We associate to the above equation the following one step Euler scheme for $X^{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ independent of $N_{\mathcal{L}(X_{s,r}^{1})}$.
$X_{s,t}^{2}(X^{2})=X^{2}+\int_{s}^{t}b(X^{2})dr+\int_{s}^{t}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,X^{2})d\widetilde{N}_{\mathcal{L}(X^{2})}(v,z,r).$ (5.34)
where $dN_{\mathcal{L}(X^{2})}(v,z,r)$ is a Poisson point measure independent of $X^{2}$ with compensator $d\widehat{N}_{\mathcal{L}(X^{2})}(v,z,r)=\mathcal{L}(X^{2})(dv)\nu(dz)dr$ and $\widetilde{N}_{\mathcal{L}(X^{2})}=N_{\mathcal{L}(X^{2})}-\widehat{N}_{\mathcal{L}(X^{2})}.$
Moreover, we define the maps for $\mathcal{L}(X)=\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$
$\theta_{s,t}(\mu)=\mathcal{L}(X_{s,t}^{1}(X)),\quad\Theta_{s,t}(\mu)=\mathcal{L}(X_{s,t}^{2}(X)).$
We will work under the following hypotheses:
$\left\langle x-y,b(x)-b(y)\right\rangle\leq$ $-\overline{b}\left|x-y\right|^{2}\quad and\quad\left|b(x)-b(y)\right|^{2}\leq C_{b}\left|x-y\right|^{2}$ (5.35)
$\left|c(v,z,x)-c(v^{\prime},z,x^{\prime})\right|^{2}\leq$ $\overline{c}^{2}(z)(\left|v-v^{\prime}\right|^{2}+\left|x-x^{\prime}\right|^{2})\quad and$ (5.36)
$\left|c(0,z,0)\right|^{2}$ $\leq\overline{c}^{2}(z)\quad with$ (5.37)
$Q:=$ $\int_{E}\overline{c}^{2}(z)\nu(dz)<\infty.$ (5.38)
Under these hypotheses, [ABC24, Theorem 3.1] gives the existence of a solution to Equation (5.33) and $\theta_{s,t}(\mu)$ is uniquely defined. Furthermore, it is also proven that
$W_{2}(\theta_{s,t}(\mu),\Theta_{t^{n}_{n},t^{n}_{n-1}}\circ\dots\circ\Theta_{t^{n}_{0},t^{n}_{1}}(\mu))\to_{n\to\infty}0$ (5.39)
for any sequence $s=t^{n}_{0}<t^{n}_{1}<\dots<t^{n}_{n}=t$ such that $\max_{1\leq i\leq n}t^{n}_{i}-t^{n}_{i-1}\to_{n\to\infty}0$.
Theorem 5.7.
Assume that (5.355.38) hold with $Q<\overline{b}$. Then, $\theta$ and $\Theta$ satisfy the $2$-Foster-Lyapunov condition, are $(2,b_{\ast},1/2)$ coupled for every $b_{\ast}<2(\bar{b}-Q)$, and $\theta$ is self-coupled with the same parameters.
Furthermore, the Markov process $(\mathcal{X}_{s,t}^{1}(X))_{t\geq s}$ admits a unique invariant measure $\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$, and for any $\beta\in(0,2(\bar{b}-Q))\cap(0,1/2]$ there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in{\mathcal{P}}_{2}({\mathbb{R}})$, $n\geq 1$ and $t>s$,
$W_{2}^{2}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq C(\left\|\mu\right\|_{2}^{2}+1)\frac{1}{n^{\beta}},\ W_{2}^{2}(\theta_{s,t}(\mu),\nu)\leq C(\left\|\mu\right\|_{2}^{2}+1)\times e^{-\beta(t-s)},$
where $t_{k}=s+\sum_{i=1}^{k}\frac{1}{1+i}$ and $\pi=\{t_{0}<\dots<t_{n}\}$.
The proof of this theorem follows the same steps as the one with $W_{1}$ estimates and is given in the Appendix B.
Appendix A Proof of Lemma 2.5
Proof of A. We first suppose that $b>\varpi(\pi)\varepsilon$ and we prove that $\sigma_{b,\varepsilon}(n)\leq C_{b,\varepsilon}\gamma_{n}^{\varepsilon}.$ Let $\eta>0$ be such that $(\varpi(\pi)+\eta)\varepsilon<b$. Notice first that by the definition of $\varpi(\pi)$, there exists $n_{0}$ such that for $n\geq n_{0}$, have
$\frac{\gamma_{n}}{\gamma_{n+1}}\leq 1+(\varpi(\pi)+\eta)\gamma_{n+1}\leq e^{(\varpi(\pi)+\eta)\gamma_{n+1}}.$ (A.1)
Since $n_{0}$ depends on the sequence $(\gamma_{n})_{n\in{\mathbb{N}}}$, so does the constant $C_{n,\varepsilon}$ that we construct now.  We prove that the sequence $v_{n}:=\sigma_{b,\varepsilon}(n)\times\gamma_{n}^{-\varepsilon}$ is bounded. We have the recurrence relation for $n\in{\mathbb{N}}$ (note that $v_{0}=0$)
$v_{n+1}=\theta_{n}v_{n}+\gamma_{n+1}\quad\text{ with }\quad\theta_{n}=\frac{\gamma_{n}^{\varepsilon}}{\gamma_{n+1}^{\varepsilon}}\times e^{-b\gamma_{n+1}}.$
Using (A.1), we get $v_{n+1}\leq e^{(\varepsilon(\varpi(\pi)+\eta)-b)\gamma_{n+1}}v_{n}+\gamma_{n+1}$ for $n\geq n_{0}$ and thus
$e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n+1}}v_{n+1}\leq e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n}}v_{n}+e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n+1}}\gamma_{n+1}.$
Using recursively this inequality, we obtain for $n>n_{0}$
$e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n}}v_{n}$ $\leq$ $e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n_{0}}}v_{n_{0}}+\sum_{i=n_{0}}^{n-1}e^{(b-\varepsilon\varpi(\pi)+\eta)t_{i+1}}\times\gamma_{i+1}$
$\leq$ $e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n_{0}}}v_{n_{0}}+e^{(b-\varepsilon(\varpi(\pi)+\eta))}\sum_{i=n_{0}}^{n-1}e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{i}}\times\gamma_{i+1}$
$\leq$ $e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n_{0}}}v_{n_{0}}+e^{(b-\varepsilon(\varpi(\pi)+\eta))}\int_{t_{n_{0}}}^{t_{n}}e^{(b-\varepsilon(\varpi(\pi)+\eta))s}ds$
$=$ $e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n_{0}}}v_{n_{0}}+e^{(b-\varepsilon(\varpi(\pi)+\eta))}\times\frac{e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n}}-e^{(b-\varepsilon(\varpi(\pi)+\eta))t_{n_{0}}}}{b-\varepsilon(\varpi(\pi)+\eta)},$
where we have used that $b-\varepsilon(\varpi(\pi)+\eta)>0$ and $\gamma_{i+1}\in(0,1)$ for $i\geq 0$. We finally get
$v_{n}\leq v_{n_{0}}+\frac{e^{(b-\varepsilon(\varpi(\pi)+\eta))}}{b-\varepsilon(\varpi(\pi)+\eta)},$
and thus the boundedness of the sequence $(v_{n})_{n\in{\mathbb{N}}}.$
Suppose now $b\leq\varpi(\pi)\varepsilon.$ We take $\lambda<\frac{b}{\varpi(\pi)}\leq\varepsilon.$ We have $b>\varpi(\pi)\lambda$ so that, in view of the previous step,  $\sigma_{b,\varepsilon}(n)\leq\sigma_{b,\lambda}(n)\leq C_{b,\lambda}\gamma_{n}^{\lambda}.$ So A is proved.
Proof of B. We now have $\gamma_{n}=\frac{1}{n+h^{-1}}.$ Clearly $\varpi(\pi)=1$ and, if $b>\varepsilon$ we use the result from the point A directly. Suppose now the $b\leq\varepsilon.$ One also has
$\ln\frac{n+h^{-1}}{k+h^{-1}}\leq\int_{k}^{n}\frac{dx}{x+h^{-1}}\leq\sum_{i=k}^{n-1}\gamma_{i}=t_{n-1}-t_{k-1}$
which gives
$e^{-b(t_{n}-t_{k})}\leq\frac{(k+1+h^{-1})^{b}}{(n+1+h^{-1})^{b}}$
and consequently
$\sigma_{b,\varepsilon}(n)=\sum_{k=1}^{n}e^{-b(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}\leq\frac{1}{(n+1+h^{-1})^{b}}\sum_{k=1}^{n}(k+1+h^{-1})^{b}\gamma_{k}^{1+\varepsilon}\leq\frac{(1+h)^{b}}{(n+1+h^{-1})^{b}}\sum_{k=1}^{n}\gamma_{k}^{1+\varepsilon-b}.$
If $b<\varepsilon$ the above quantity is bounded by $Cn^{-b}$ and if $b=\varepsilon$, it is upper bounded by $Cn^{-b}\ln(n+1)$. This gives the first upper bound.
We have
$t_{n}-s=\sum_{i=1}^{n}\frac{1}{i+h^{-1}}\leq\int_{h^{-1}}^{n+h^{-1}}\frac{dx}{x}=\ln(n+h^{-1})-\ln h^{-1}$
so that
$he^{-(t_{n}-s)}\geq\frac{1}{n+h^{-1}}$ (A.2)
which also gives, for $n\geq h^{-1},$
$\frac{1}{n}=\frac{1}{n+h^{-1}}\times\frac{n+h^{-1}}{n}\leq 2he^{-(t_{n}-s)}\leq 2ehe^{-(t-s)},$
since $0\leq t-t_{n}\leq t_{n+1}-t_{n}\leq 1$. Then, if $b\neq\varepsilon$ one has
$\sigma_{b,\varepsilon}(n)\leq C_{b,\varepsilon}(\frac{2e}{h})^{b\wedge\varepsilon}e^{-b\wedge\varepsilon(t-s)}.$
If $b=\varepsilon$ we have to multiply with $\ln(1+n)\leq C^{\prime}(1+(t-s))$. $\square$
Appendix B Additional technical proofs for some applications
B.1. A stochastic equation with reflection in $W_{2}$: Proof of Theorem 5.1
We check the conditions in Theorem 2.15 in two situations: on the one hand to get that $\mathcal{X}^{1}$ is self-coupled, and on the other hand to obtain that $\mathcal{X}^{1}$ is coupled with $\mathcal{X}^{2}$, where $\mathcal{X}^{1}$ and $\mathcal{X}^{2}$ are defined respectively by (5.1) and (5.4). In the present case, the coupling is the simplest one, which is given by the strong solution of the corresponding reflected equation driven by the same Brownian motion.
Furthermore, note that from the contraction property (5.5), we obtain that for every $\delta>0$,
$xb(x)\leq$ $(-\bar{b}+\delta)|x|^{2}+\delta^{-1}b(0)^{2}.$ (B.1)
First, we give a lemma with some moment estimates.
Lemma B.1.
Assume that $2\bar{b}>C_{\sigma}$, then there exists a positive constant $C$ that depends on $C_{b}$, $C_{\sigma}$, $b(0)$ and $\sigma(0)$ such that for $s\leq t\leq s+1$
${\mathbb{E}}[(\mathcal{X}_{s,t}^{i}(X^{i})-X^{i})^{2}]\leq$ $C(1+{\mathbb{E}}[(X^{i})^{2}])(t-s),$ (B.2)
$\left|{\mathbb{E}}(\mathcal{X}_{s,t}^{i}(X^{i})^{2})-{\mathbb{E}}[(X^{i})^{2}]\right|$ $\leq C(1+{\mathbb{E}}[(X^{i})^{2}])(t-s).$ (B.3)
Besides, (3.4) is satisfied with $p=2$, $\varepsilon^{\prime}=1$ for $\mathcal{X}_{s,t}^{i}(X^{i})$ and $\mathcal{X}_{s,t}^{1}(X^{i})$, $i=1,2$ .
Proof.
Using Itô’s formula and the fact that $\mathcal{X}_{s,u}^{i}(X^{i})dL^{i}_{s,u}=0$ for $i\in\{1,2\}$, we obtain
${\mathbb{E}}[\mathcal{X}_{s,t}^{1}(X^{1})^{2}]$ ${=}{\mathbb{E}}[(X^{1})^{2}]+2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}_{s,u}^{1}(X^{1})b(\mathcal{X}_{s,u}^{1}(X^{1}))]du+\int_{s}^{t}{\mathbb{E}}[\sigma(\mathcal{X}_{s,u}^{1}(X^{1}))^{2}]du$ (B.4)
${\mathbb{E}}[\mathcal{X}_{s,t}^{2}(X^{2})^{2}]$ ${=}{\mathbb{E}}[(X^{2})^{2}]+2\int_{s}^{t}{\mathbb{E}}[\mathcal{X}_{s,u}^{2}(X^{2})b(X^{2})]du+\int_{s}^{t}{\mathbb{E}}[\sigma(X^{2})^{2}]du.$ (B.5)
Using the inequality $2ab\leq a^{2}+b^{2}$ then the Lipschitz properties in (5.3) together with Gronwall inequality, we get that there exists $C$ depending on $C_{b}$, $C_{\sigma}$, $|b(0)|$ and $|\sigma(0)|$ such that
${\mathbb{E}}[\mathcal{X}_{s,t}^{i}(X^{i})^{2}]\leq C(1+{\mathbb{E}}[(X^{i})^{2}]),\ s\leq t\leq s+1,\ i\in\{1,2\}.$ (B.6)
In the same way, we get from Itô’s formula, the inequality $2ab\leq a^{2}+b^{2}$, $(\mathcal{X}_{s,t}^{i}(X^{i})-X^{i})dL_{s,t}^{i}=-X^{i}dL_{s,t}^{i}\leq 0$ and (B.6):
${\mathbb{E}}[(\mathcal{X}_{s,t}^{i}(X^{i})-X^{i})^{2}]{\leq}\int_{s}^{t}{\mathbb{E}}[(\mathcal{X}_{s,u}^{i}(X^{i})-X^{i})^{2}]du+C{\mathbb{E}}(1+|X^{i}|^{2})(t-s),$
where $C$ depends on $C_{b}$, $C_{\sigma}$, $|b(0)|$ and $|\sigma(0)|$. Applying Gronwall’s inequality one obtains the preliminary estimate (B.2). In a similar fashion, we get (B.3). ∎
Lemma B.2.
$\theta$ and $\Theta$ satisfy the $2$-Foster-Lyapunov condition.
Proof.
We will prove the property (3.4) which is equivalent to (3.2). From (B.4), (B.1) and $\sigma(x)^{2}\leq(1+\delta)C_{\sigma}|x|^{2}+(1+\delta^{-1})\sigma(0)^{2}$, we get
${\mathbb{E}}[\mathcal{X}_{s,t}^{1}(X^{1})^{2}]$ ${\leq}$ ${\mathbb{E}}[(X^{1})^{2}]-(2(\bar{b}-\delta)-C_{\sigma}(1+\delta))\int_{s}^{t}{\mathbb{E}}(\mathcal{X}_{s,u}^{1}(X^{1})^{2})du$
$+\left(\delta^{-1}(\sigma^{2}(0)+2b^{2}(0))+\sigma^{2}(0)\right)(t-s)$
${=}$ ${\mathbb{E}}[(X^{1})^{2}](1-(2\bar{b}-C_{\sigma}-(C_{\sigma}+2)\delta)(t-s))$
$-(2\bar{b}-C_{\sigma}-(C_{\sigma}+2)\delta)\int_{s}^{t}({\mathbb{E}}(\mathcal{X}_{s,u}^{1}(X^{1}))^{2}-{\mathbb{E}}(X^{1})^{2})du$
$+\left(\delta^{-1}(\sigma^{2}(0)+2b^{2}(0))+\sigma^{2}(0)\right)(t-s)$
$\leq$ ${\mathbb{E}}[(X^{1})^{2}](1-(2\bar{b}-C_{\sigma}-(C_{\sigma}+2)\delta)(t-s))+C(1+{\mathbb{E}}[(X^{1})^{2}])(t-s)^{2}$
$+\left(\delta^{-1}(\sigma^{2}(0)+2b^{2}(0))+\sigma^{2}(0)\right)(t-s),$
by using (B.3) for the last inequality. Now take $\delta>0$ sufficiently small so that $2\bar{b}-C_{\sigma}-(C_{\sigma}+2)\delta>0$. This is precisely (3.4) for $p=2$ applied to $\theta$ with $\varepsilon=1$.
A small variation of the above argument gives the result for ${\mathbb{E}}[\mathcal{X}_{s,t}^{2}(X^{2})^{2}]$. In fact, using Itô’s formula and $\mathcal{X}_{s,u}^{2}(X^{2})dL^{2}_{s,u}=0$, we obtain using (B.1)
${\mathbb{E}}[\mathcal{X}_{s,t}^{2}(X^{2})^{2}]{=}$ ${\mathbb{E}}[(X^{2})^{2}]+2\int_{s}^{t}{\mathbb{E}}[(\mathcal{X}_{s,u}^{2}(X^{2})-X^{2})b(X^{2})]du+2{\mathbb{E}}[X^{2}b(X^{2})](t-s)$
$+(t-s){\mathbb{E}}[\sigma(X^{2})^{2}]$
$\leq$ ${\mathbb{E}}[(X^{2})^{2}](1-2(\bar{b}-\delta)(t-s))+(t-s){\mathbb{E}}[\sigma(X^{2})^{2}+\delta b(X^{2})^{2}]$
$+\delta^{-1}\int_{s}^{t}{\mathbb{E}}[(\mathcal{X}_{s,u}^{2}(X^{2})-X^{2})^{2}]du+\delta^{-1}b(0)^{2}(t-s)$
$\leq$ ${\mathbb{E}}[(X^{2})^{2}](1-(2\bar{b}-(2+2C_{b}+C_{\sigma})\delta-C_{\sigma})(t-s))+\delta^{-1}C(t-s)^{2}(1+{\mathbb{E}}[(X^{2})^{2}]),$
by using (B.2) and the Lipschitz property of $b$ and $\sigma$. Taking $\delta>0$ sufficiently small, we thus get (3.4) with $\varepsilon=1$. ∎
The next lemma shows that the condition (2.22) is satisfied.
Lemma B.3.
Suppose that $2\bar{b}-C_{\sigma}>0$ then (2.22) is satisfied for $p=2$, $\varepsilon=1$ and any $b_{*}\in(0,2\bar{b}-C_{\sigma})$ for the following two pairs $(\mathcal{X}_{s,t}^{1}(X^{1}),\mathcal{X}_{s,t}^{2}(X^{2}))$ and $(\mathcal{X}_{s,t}^{1}(X^{1}),\mathcal{X}_{s,t}^{1}(X^{2}))$. The constant $C_{*}$ in (2.22) depends on $\bar{b}$, $C_{b}$, $C_{\sigma}$, $b(0)$ and $\sigma(0)$.
Proof.
We proceed in a similar fashion as in the proof of Lemma B.1. Define $\mathcal{Y}:=\mathcal{X}^{1}(X^{1})-\mathcal{X}^{2}(X^{2})$ which satisfies, for $s\leq t$,
$\mathcal{Y}_{s,t}=X^{1}-X^{2}$ $+\int_{s}^{t}\left(b(\mathcal{X}_{s,r}^{1}(X^{1}))-b(X^{2})\right)dr$
$+\int_{s}^{t}\left(\sigma(\mathcal{X}_{s,r}^{1}(X^{1}))-\sigma(X^{2})\right)dW_{r}+L^{1}_{s,t}-L^{2}_{s,t}.$
First, we prove the following preliminary estimates:
$\left|{\mathbb{E}}[\mathcal{Y}_{s,t}^{2}]-{\mathbb{E}}[(X^{1}-X^{2})^{2}]\right|$ $\leq C(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(t-s),$ (B.7)
$\left|{\mathbb{E}}[(\mathcal{X}_{s,t}^{1}(X^{1})-\mathcal{X}_{s,t}^{1}(X^{2}))^{2}]-{\mathbb{E}}[(X^{1}-X^{2})^{2}]\right|$ $\leq C(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(t-s).$ (B.8)
where the constant $C$ depends on $C_{b}$ and $C_{\sigma}$ as well as $b(0)$ and $\sigma(0)$. We give the argument for (B.7), leaving (B.8) for the reader. Using Itô’s formula and $(\mathcal{X}_{s,u}^{1}(X^{1})-\mathcal{X}_{s,u}^{2}(X^{2}))(dL_{s,u}^{1}-dL_{s,u}^{2})\leq 0$, we obtain for $t-s\leq 1$
${\mathbb{E}}[(\mathcal{Y}_{s,t})^{2}]{\leq}$ ${\mathbb{E}}[(X^{1}-X^{2})^{2}]+2\int_{s}^{t}{\mathbb{E}}[\mathcal{Y}_{s,u}(b(\mathcal{X}_{s,u}^{1}(X^{1}))-b(X^{2}))]du$
$+\int_{s}^{t}{\mathbb{E}}[(\sigma(\mathcal{X}_{s,u}^{1}(X^{1}))-\sigma(X^{2}))^{2}]du.$ (B.9)
Therefore, using the the inequality $2ab\leq a^{2}+b^{2}$ for the second term and then Lipschitz properties of $b$ and $\sigma$ as well as bounds for the $L^{2}$-moments of $\mathcal{X}^{i}(X^{i})$, $i=1,2$ in Lemma B.1 , (B.7) follows.
In order to prove (2.22), we use again (B.9) but we apply the inequality $ab\leq\delta a^{2}+\delta^{-1}b^{2}$, the Lipschitz property of $b$ and the following decomposition using (5.5)
$2{\mathbb{E}}[\mathcal{Y}_{s,u}$ $(b(\mathcal{X}_{s,u}^{1}(X^{1}))-b(X^{2}))]\leq-2\bar{b}{\mathbb{E}}[\mathcal{Y}_{s,u}^{2}]+2{\mathbb{E}}[\mathcal{Y}_{s,u}(b(\mathcal{X}_{s,u}^{2}(X^{2}))-b(X^{2}))]$
$\leq$ $-(2\bar{b}-\delta){\mathbb{E}}[\mathcal{Y}_{s,u}^{2}]+C_{b}\delta^{-1}{\mathbb{E}}[(\mathcal{X}_{s,u}^{2}(X^{2})-X^{2})^{2}]$
$\leq$ $-(2\bar{b}-\delta){\mathbb{E}}[(X^{1}-X^{2})^{2}]+C(\delta+\delta^{-1})(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(u-s).$
In the last inequality, we have used (B.7) for the first term, and Lemma B.1 for the second one. Similarly, using the inequality $ab\leq\delta a^{2}+\delta^{-1}b^{2}$ with ${\mathbb{E}}[(\sigma(\mathcal{X}_{s,u}^{2}(X^{2}))-\sigma(X^{2}))^{2}]\leq C(1+{\mathbb{E}}[(X^{2})^{2}])(u-s)$ by (B.2), we get
${\mathbb{E}}[(\sigma(\mathcal{X}_{s,u}^{1}(X^{1}))-\sigma(X^{2}))^{2}]\leq$ $C_{\sigma}(1+\delta){\mathbb{E}}[\mathcal{Y}^{2}_{s,u}]+(1+\delta^{-1})C(1+{\mathbb{E}}[(X^{2})^{2}])(u-s)$
$\leq$ $C_{\sigma}(1+\delta){\mathbb{E}}[(X^{1}-X^{2})^{2}]$
$+(1+\delta^{-1})C(1+{\mathbb{E}}[(X^{1})^{2}]+{\mathbb{E}}[(X^{2})^{2}])(u-s).$
Therefore putting the above estimates together and taking $\delta>0$ sufficiently small, we obtain (2.22) for $\varepsilon=1$ and any $b_{*}\in(0,2\bar{b}-C_{\sigma})$. Similar arguments give (2.22) for $\mathcal{X}_{s,t}^{1}(X^{i})$, $i=1,2$. We conclude by applying Theorem 2.12: since $b_{*}$ can take any value in $(0,2\bar{b}-C_{\sigma})$, we take $b_{*}\in(1,2\bar{b}-C_{\sigma})$ if $2\bar{b}-C_{\sigma}>1$ to get the convergence with rate one and $\beta=b_{*}$ otherwise. ∎
B.2. A neuronal model in $W_{1}$: Proof of Theorem 5.4
For $0\leq t-s\leq 1$, we obtain by taking expectations on the stochastic equation that defines $\mathcal{X}^{i}_{s,\cdot}(X^{i})$ in (5.8), we have for $i=1,2$ (recall that $\mathcal{X}^{i}_{s,t}(X^{i})\geq 0$),
${\mathbb{E}}[\mathcal{X}^{i}_{s,t}(X^{i})]-{\mathbb{E}}[X^{i}]=$ $\int_{s}^{t}{\mathbb{E}}[b_{i}(r)+J{\mathbb{E}}[f_{i}(r)]]dr-\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{i}_{s,r}(X^{i})f_{M}(\mathcal{X}^{i}_{s,r}(X^{i}))]dr.$ (B.10)
Next using the Lipschitz property (5.7)
${\mathbb{E}}[\mathcal{X}^{i}_{s,t}(X^{i})]\leq{\mathbb{E}}(X^{i})+b(0)(t-s)+(C_{b}+JC_{f})\int_{s}^{t}{\mathbb{E}}[\mathcal{X}^{i}_{s,u}]du,$
since the jump contribution is nonpositive. This gives
${\mathbb{E}}(\mathcal{X}^{i}_{s,t}(X^{i}))\leq C(1+{\mathbb{E}}(X^{i})),$ (B.11)
with $C$ depending on $b(0)$, $C_{f}$ and $J$.
With these estimates we can now prove (3.4). We use Itô’s fomula for $F(y)=|y|$, and we use a regularization procedure as in the proof of Theorem 5.6. We get
${\mathbb{E}}[|\mathcal{X}^{i}_{s,t}(X^{i})-X^{i}|]$ $=\int_{s}^{t}{\mathbb{E}}[1_{\mathcal{X}^{i}_{s,r}(X^{i})-X^{i}>0}(b_{i}(r)+J{\mathbb{E}}[f_{i}(r)])]dr$
$+\int_{s}^{t}{\mathbb{E}}[(|X^{i}|-|\mathcal{X}^{i}_{s,r}(X^{i})-X^{i}|)f_{M}(\mathcal{X}^{i}_{s,r}(X^{i}))]dr.$
Using (B.11), we obtain for $i=1$
${\mathbb{E}}[|\mathcal{X}^{1}_{s,t}(X^{1})-X^{1}|]$ $\leq\int_{s}^{t}b(0)+(C_{b}+C_{f}J+M){\mathbb{E}}[\mathcal{X}^{1}_{s,r}(X^{1})]dr$
$\leq C(1+{\mathbb{E}}[X^{1}])(t-s).$ (B.12)
For $i=2$ the estimation is done similarly. Note that these two inequalities also imply that for $\mathcal{Y}_{s,t}:=\mathcal{X}^{1}_{s,t}(X^{1})-\mathcal{X}^{2}_{s,t}(X^{2})$, we have
$\left|{\mathbb{E}}[|\mathcal{Y}_{s,t}|-|\mathcal{Y}_{s,s}|]\right|\leq C(1+{\mathbb{E}}[X^{1}]+{\mathbb{E}}[X^{2}])(t-s).$ (B.13)
The contraction hypothesis (5.17) with $x=0$ gives $b(y)-yf_{M}(y)\leq b(0)-\bar{b}y$ for $y\geq 0$, and we thus obtain applying this inequality to (B.10)
${\mathbb{E}}(\mathcal{X}^{i}_{s,t}(X^{i}))-{\mathbb{E}}(X^{i})\leq\int_{s}^{t}(-\bar{b}+JC_{f}){\mathbb{E}}[\mathcal{X}^{i}_{s,r}(X^{i})]dr+b(0)(t-s)+C(1+{\mathbb{E}}(X^{i}))(t-s)^{2},$
by using (B.12) for $i=2$. We now use that $\mathcal{X}^{i}_{s,r}(X^{i})\leq X^{i}+|\mathcal{X}^{i}_{s,r}(X^{i})-X^{i}|$ with (B.12) to obtain (3.4) for $p=1$ with $\varepsilon=1$.
Next, in order to prove (2.22), we write
$\mathcal{Y}_{s,t}=\mathcal{Y}_{s,s}+\int_{s}^{t}b(\mathcal{X}^{1}_{s,r}(X^{1}))-b(X^{2})dr+J\int_{s}^{t}{\mathbb{E}}[f(\mathcal{X}^{1}_{s,r}(X^{1}))]-{\mathbb{E}}[f({X}^{2})]ds-\int_{s}^{t}\int_{\mathbb{R}_{+}}c(r,z)dN(r,z).$
with
$c(r,z)=\mathcal{X}^{1}_{s,r}(X^{1})1_{z\leq f_{M}(\mathcal{X}^{1}_{s,r}(X^{1}))}-\mathcal{X}^{2}_{s,r}(X^{2})1_{z\leq f_{M}(\mathcal{X}^{2}_{s,r}(X^{2}))}$
Then, regularizing $y\mapsto|y|$ and using Itô’s formula and taking the limit as $\varepsilon\rightarrow 0$ we get
$|\mathcal{Y}_{s,t}|=$ $|\mathcal{Y}_{s,s}|+\int_{s}^{t}\mathrm{sgn}(\mathcal{Y}_{s,r})(b(\mathcal{X}^{1}_{s,r}(X^{1}))-b({X}^{2}))dr$
$+J\int_{s}^{t}\mathrm{sgn}(\mathcal{Y}_{s,r})({\mathbb{E}}[f(\mathcal{X}^{1}_{s,r}(X^{1}))]-{\mathbb{E}}[f(X^{2})])dr$
$+\int_{s}^{t}\int_{\mathbb{R}_{+}}|\mathcal{Y}_{s,r-}-c(r,z)|-|\mathcal{Y}_{s,r-}|dN(r,z)$
$=:$ $|\mathcal{Y}_{s,s}|+\sum_{i=1}^{3}I_{i}(s;t).$
We consider the estimate for ${\mathbb{E}}[I_{3}(s;t)]$. We have
${\mathbb{E}}[I_{3}(s;t)]=$ ${\mathbb{E}}\int_{s}^{t}\big{(}-|\mathcal{Y}_{s,r}|f_{M}(\mathcal{X}^{1}_{s,r}(X^{1}))\wedge f_{M}(\mathcal{X}^{2}_{s,r}(X^{2}))$
$+(\mathcal{X}^{2}_{s,r}(X^{2})-|\mathcal{Y}_{s,r}|)(f_{M}(\mathcal{X}^{1}_{s,r}(X^{1}))-f_{M}(\mathcal{X}^{2}_{s,r}(X^{2})))^{+}$
$+(\mathcal{X}^{1}_{s,r}(X^{1})-|\mathcal{Y}_{s,r}|)(f_{M}(\mathcal{X}^{2}_{s,r}(X^{2}))-f_{M}(\mathcal{X}^{1}_{s,r}(X^{1})))^{+}\big{)}\ dr.$
As in the $W_{2}$ case, we denote by $(y,x)$ the variables $(\mathcal{X}^{1}_{s,u}(X^{1}),\mathcal{X}^{2}_{s,u}(X^{2}))$ and write
$-|y-x|f_{M}(x)\wedge f_{M}(y)+(x-|y-x|)(f_{M}(y)-f_{M}(x))^{+}+(y-|y-x|)(f_{M}(x)-f_{M}(y))^{+}$
$=$ $-|y-x|(f_{M}(y)\vee f_{M}(x))+x(f_{M}(y)-f_{M}(x))^{+}+y(f_{M}(x)-f_{M}(y))^{+}.$
Therefore, using the contraction hypothesis (5.17) and the Lipschitz property (5.7), we get
${\mathbb{E}}|\mathcal{Y}_{s,t}|$ $\leq{\mathbb{E}}|\mathcal{Y}_{s,s}|+{\mathbb{E}}\left[\int_{s}^{t}(-\bar{b}+JC_{f})|\mathcal{Y}_{s,r}|dr\right]+(C_{b}+C_{f}J){\mathbb{E}}\int_{s}^{t}|\mathcal{X}^{2}_{s,r}(X^{2})-X^{2}|dr$
$\leq{\mathbb{E}}|\mathcal{Y}_{s,s}|+{\mathbb{E}}\left[\int_{s}^{t}(-\bar{b}+JC_{f})|\mathcal{Y}_{s,r}|dr\right]+C(1+{\mathbb{E}}|X^{2}|)(t-s)^{2},$
by using (B.12) for the last inequality. Using (B.13), we see that (2.22) is satisfied for $b_{*}=\bar{b}-JC_{f}$ since $e^{-(\bar{b}-JC_{f})t}=_{t\to 0}1-(\bar{b}-JC_{f})t+O(t^{2})$.
The proof that $\theta$ is self-coupled can be obtained as in the above proof.
B.3. A Boltzmann equation in the martingale regime ($W_{2}$): Proof of Theorem 5.7
In contrast with the $W_{1}$ case (see Theorem 5.6), we are not able to deduce the coupling property from the sewing lemma, see Eq. (3.1). Thus, we prove here directly the coupling property. To construct the coupling, we work with approximation schemes on a regular time grid and let then the time step go to zero to obtain the coupling property for the flow. Namely, for each $m\in{\mathbb{N}}$ we consider the grid
$s_{i}=s+\frac{i}{m},\quad i\in{\mathbb{N}},$
and we define
$\Theta_{s,s_{i}}^{m}(\mu)=\Theta_{s_{i-1},s_{i}}\circ....\circ\Theta_{s_{0},s_{1}}(\mu)\text{ and }\Theta_{s,t}^{m}(\mu)=\Theta_{s_{i},t}\circ\Theta_{s,s_{i}}^{m}(\mu)\text{ for }t\in(s_{i},s_{i+1}).$
In order to prove Theorem 5.7, we use Theorem 2.15 so we have to produce a probabilistic representation and to prove that this representation verifies (2.22) and (3.4). For the moment we fix $m$, and for $s_{i-1}\leq u<s_{i}$ we denote $\eta(u)=s_{i-1}$.
Step 1: construction of the coupling process (stochastic representation) for $\Theta_{s,s_{i}}^{m}(\mu)$ and $\Theta_{s,s_{i}}(\mu).$ Our first step is to produce the joint representation (coupling process) for them. Let $\mu^{1},\mu^{2}$ be fixed. For each $i\in{\mathbb{N}}$, we define $\Pi_{s,s_{i}}(dv_{1},dv_{2})$ as the optimal $W_{2}$-coupling of $\Theta_{s,s_{i}}^{m}(\mu^{1})$ and $\mu^{2},$ and we take $\tau_{s,s_{i}}=(\tau_{s,s_{i}}^{1},\tau_{s,s_{i}}^{2}):(0,1)\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}F(v_{1},v_{2})\Pi_{s,s_{i}}(dv_{1},dv_{2})=\int_{0}^{1}F(\tau_{s,s_{i}}(w))dw.$
The function $\tau$ is constructed such that $(s,w)\mapsto\tau_{s,s_{i}}(w)$ is measurable, see Carmona and Delarue [CD18, Lemma 5.29].
We construct now a joint probabilistic representation (in the sense defined in Section 1). We consider a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ that endows $dN(w,z,r)$ a Poisson point measure of intensity $dw\nu(dz)dr$. We denote by $(\mathcal{F}_{t})_{t\geq 0}$ the filtration generated by this random measure and $d\tilde{N}$ the compensated measure. Moreover, we consider random variables $X^{i}\in L^{2}(\Omega),i=1,2$ which are $\mathcal{F}_{s}$ measurable and such that $(X^{1},X^{2})$ has the law $\Pi_{s,s}(dv_{1},dv_{2}),$ which is the $W_{2}$ optimal coupling of $\mu^{1}$ and $\mu^{2}.$ In particular $X^{i}\sim\mu^{i}$. Then, we define the equations which give the coupling
$x_{s,t}^{1}(X^{1})$ $=X^{1}+\int_{s}^{t}b(x_{s,\eta(r)}^{1}(X^{1}))dr+\int_{s}^{t}\int_{E}\int_{0}^{1}c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1}(X^{1}))d\widetilde{N}(w,z,r)$
$x_{s,t}^{2}(X^{2})$ $=X^{2}+\int_{s}^{t}b(X^{2})dr+\int_{s}^{t}\int_{E}\int_{0}^{1}c(\tau_{s,\eta(r)}^{2}(w),z,X^{2})d\widetilde{N}(w,z,r)$ (B.14)
We remark that by induction all above integrals driven by $\widetilde{N}$ are square integrable martingales and therefore their expectation is zero. This and similar arguments for other integrals driven by $\widetilde{N}$ will be frequently used in what follows.
Lemma B.4.
For any $t\geq s$, $x_{s,t}^{1}(X^{1})$ has the law $\Theta_{s,t}^{m}(\mu^{1}),$ and $x_{s,t}^{2}(X^{2})$ has the law $\Theta_{s,t}(\mu^{2})$.
Proof.
Notice first that $\Theta_{s,t}^{m}(\mu^{1})$ is the law of $X_{s,t}^{1}(X^{1})$ constructed recursively from (5.34) by
$X_{s,t}^{1}(X^{1})$ $=$ $X^{1}+\int_{s}^{t}b(x_{s,\eta(r)}^{1}(X^{1}))dr+\sum_{s_{j}\leq t}\int_{s_{j-1}}^{s_{j}}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,X_{s,s_{j-1}}^{1}(X^{1}))d\widetilde{N}_{\rho_{r}}(v,z,r)$
$+\int_{\eta(t)}^{t}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,x_{s,\eta(t)}^{1}(X^{1}))d\widetilde{N}_{\rho_{r}}(v,z,r)$
with $\rho_{r}(dv)=\Theta_{s,s_{j-1}}^{m}(\mu^{1})(dv)$ for $s_{j-1}\leq r<s_{j}.$ Let
$I_{j}(x)$ $=$ $\int_{s_{j-1}}^{s_{j}}\int_{E}\int_{0}^{1}c(\tau_{s,s_{j-1}}^{1}(w),z,x)d\widetilde{N}(w,z,r)$
$I_{j}^{\prime}(x)$ $=$ $\int_{s_{j-1}}^{s_{j}}\int_{E}\int_{\mathbb{R}^{d}}c(v,z,x)d\widetilde{N}_{\rho_{r}}(v,z,r).$
Using Itô’s formula one easily checks that
${\mathbb{E}}(e^{i\left\langle h,I_{j}(x)\right\rangle})=e^{(s_{j}-s_{j-1})\lambda_{h}(x)}\quad\text{and}\quad{\mathbb{E}}(e^{i\left\langle h,I_{j}^{\prime}(x)\right\rangle})=e^{(s_{j}-s_{j-1})\lambda_{h}^{\prime}(x)}$
with
$\lambda_{h}(x)$ $=$ $\int_{0}^{1}\int_{E}\left(e^{i\left\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\right\rangle}-1-i\langle h,c(\tau_{s,s_{j-1}}^{1}(w),z,x)\rangle\right)\nu(dz)dw$
$\lambda_{h}^{\prime}(x)$ $=$ $\int_{\mathbb{R}^{d}}\int_{E}\left(e^{i\langle h,c(v,z,x)\rangle}-1-i\langle h,c(v,z,x)\rangle\right)\nu(dz)\Theta_{s,s_{j-1}}^{m}(\mu^{1})(dv).$
By the very definition of $\tau_{s,s_{j-1}}^{1}(w)$, the image of the uniform distribution is $\Theta_{s,s_{j-1}}^{m}(\mu^{1})$. Therefore, $\lambda_{h}(x)=\lambda_{h}^{\prime}(x)$ and the law of $I_{i}(x)$ coincides with the law of $I_{i}^{\prime}(x).$ Then, conditioning on $\mathcal{F}_{s_{j}}$, one proves by recurrence that $x_{s,t}^{1}(X^{1})$ has law $\Theta_{s,t}^{m}(\mu^{1})$ for $t\in[s_{j},t_{j+1}]$ and thus for any $t\geq s$. In a similar way (but without recurrence) one checks that $x_{s,t}^{2}(X^{1})$ has law $\Theta_{s,t}(\mu^{2})$. ∎
Both $x_{s,t}^{1}(X^{1})$ and $x_{s,t}^{2}(X^{2})$ are defined on the same probability space - so they represent a joint probabilistic representation. But we have to check the hypothesis concerning the coupling: see (2.21), (2.22 ),(3.4). By the very construction (2.21) is verified.
Step 2: The Foster-Lyapunov condition (3.4).
From the assumptions (5.35) and (5.36), we have for every $\delta>0$
$\left\langle x,b(x)\right\rangle$ $\leq$ $(-\overline{b}+\delta)\left|x\right|^{2}+\delta^{-1}\left|b(0)\right|^{2}$ (B.15)
$\left|c(v,z,x)\right|^{2}$ $\leq$ $\overline{c}^{2}(z)(1+\left|v\right|^{2}+\left|x\right|^{2})\quad and\quad\left|b(x)\right|\leq K(1+\left|x\right|).$ (B.16)
Lemma B.5.
Under hypotheses (5.35) and (5.36), we have the following:
A There exists a constant $C$ (depending on $Q,\overline{b}$ and $K$, but not on the time step $1/m$) such that
${\mathbb{E}}\left(|x_{s,t}^{i}(X^{i})-X^{i}|^{2}\right)\leq C(1+{\mathbb{E}}\left(|X^{i}|^{2}\right))(t-s).$ (B.17)
B If $\overline{b}>Q+\delta$ then for every $s<t\leq s+1$
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)$ $\leq$ ${\mathbb{E}}\left(|X^{1}|^{2}\right)(1-2(\overline{b}-Q-\delta)(t-s))$
$+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s)+C(1+{\mathbb{E}}\left(|X^{1}|^{2}\right))(t-s)^{3/2}.$
Note that the result in (B.5) differs in rate from the one in Lemma B.1.
Proof.
Through the proof, we simplify the notation using $(x_{s,t}^{1},x_{s,t}^{2})=(x_{s,t}^{1}(X^{1}),x_{s,t}^{2}(X^{2}))$.
We start from the equation satisfied by $x_{s,t}^{1}$ in (B.14). Using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (5.38), (B.15) and (B.16), one gets
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)\leq C\left(1+{\mathbb{E}}\left|X^{1}\right|^{2}+\int_{s}^{t}{\mathbb{E}}\left(|x_{s,\eta(r)}^{1}|^{2}\right)dr\right).$
Here, we have used that $\int_{0}^{1}|\tau_{s,\eta(r)}^{1}(w)|^{2}dw={\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})$ and the constant $C$ depends on $Q$. This identity will be used frequently without any further mention of it. Next, due to a Gronwall type inequality we obtain
${\mathbb{E}}(|x_{s,t}^{1}|^{2})\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2}).$ (B.19)
Proof of A: Let us prove (B.17). As before, note that using the inequality $|x+y|^{2}\leq 2(|x|^{2}+|y|^{2})$, Cauchy-Schwarz inequality, the $L^{2}$-isometry for martingale driven stochastic integrals, (B.16) and (B.19) one obtains
${\mathbb{E}}\left(|x_{s,t}^{1}(X^{1})-X^{1}|^{2}\right)$
$\leq$ $2(t-s){\mathbb{E}}\int_{s}^{t}|b(x_{s,\eta(r)}^{1})|^{2}dr+2{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})|^{2}dwd\nu(z)dr$
$\leq$ $2K(t-s){\mathbb{E}}\int_{s}^{t}(1+|x_{s,\eta(r)}^{1}|^{2})dr+2Q\int_{s}^{t}(1+2{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2}))dr$
$\leq$ $C^{\prime}(t-s)(1+{\mathbb{E}}\left(|X^{1}|^{2}\right)).$
The proof in the case $i=2$ is the same.
Proof of B: One employs (B.17) and (B.15) to obtain
${\mathbb{E}}\left(\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\rangle\right)$ $\leq$ $\left({\mathbb{E}}(|x_{s,r}^{1}-x_{s,\eta(r)}^{1}|^{2}){\mathbb{E}}(|b(x_{s,\eta(r)}^{1})|^{2})\right)^{1/2}+{\mathbb{E}}\left(\langle x_{s,\eta(r)}^{1},b(x_{s,\eta(r)}^{1})\rangle\right)$
$\leq$ $C(r-\eta(r))^{1/2}(1+{\mathbb{E}}(|X^{1}|^{2}))+(-\overline{b}+\delta){\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})+\delta^{-1}|b(0)|^{2}$
Since $\int_{0}^{1}\left|\tau_{s,\eta(r)}^{1}(w)\right|^{2}dw={\mathbb{E}}\left|x_{s,\eta(r)}^{1}\right|^{2}$ and we have assumed ( 5.38) we also have
${\mathbb{E}}\left(\int_{s}^{t}\int_{E}\int_{0}^{1}\overline{c}^{2}(z)(1+|\tau_{s,\eta(r)}^{1}(w)|^{2}+|x_{s,\eta(r)}^{1}|^{2})dwd\nu(z)dr\right)\leq Q\int_{s}^{t}(1+2{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2}))dr$
Then, using Itô’s formula for the function $x\rightarrow\left|x\right|^{2},$ and the identity $\left|x+c\right|^{2}-\left|x\right|^{2}-2\left\langle x,c\right\rangle=\left|c\right|^{2},$ we obtain
${\mathbb{E}}\left(|x_{s,t}^{1}|^{2}\right)=$ ${\mathbb{E}}(|X^{1}|^{2})+2{\mathbb{E}}\int_{s}^{t}\left\langle x_{s,r}^{1},b(x_{s,\eta(r)}^{1})\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})\right|^{2}dwd\nu(z)dr.$
$\leq$ ${\mathbb{E}}\left(|X^{1}|^{2}\right)+C(t-s)^{3/2}(1+{\mathbb{E}}(|X^{1}|^{2}))-2(\overline{b}-Q-\delta){\mathbb{E}}\int_{s}^{t}|x_{s,\eta(r)}^{1}|^{2}dr$
$+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s)$
We have $|{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})-{\mathbb{E}}(|X^{1}|^{2})|\leq\sqrt{{\mathbb{E}}((|x_{s,\eta(r)}^{1}|+|X^{1}|)^{2})}\sqrt{{\mathbb{E}}(|x_{s,\eta(r)}^{1}-X^{1}|^{2})}$, and we get by using A and (B.19)
$\int_{s}^{t}\left|{\mathbb{E}}(|x_{s,\eta(r)}^{1}|^{2})-{\mathbb{E}}(|X^{1}|^{2})\right|dr\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2})(t-s)^{3/2}.$
The above inequalities give
${\mathbb{E}}(|x_{s,t}^{1}|^{2})$ $\leq$ $(1-2(\overline{b}-Q-\delta)(t-s)){\mathbb{E}}\left|X^{1}\right|^{2}$
$+C(t-s)^{3/2}(1+{\mathbb{E}}(|X^{1}|^{2}))+(2\delta^{-1}\left|b(0)\right|^{2}+Q)(t-s).\qed$
We conclude from Lemma B.5 that
$\left\|\Theta_{s,t}^{m}(\mu)\right\|_{2}^{2}\leq(1-2(\overline{b}-Q-\delta))(t-s))\left\|\mu\right\|_{2}^{2}+C(1+\left\|\mu\right\|_{2}^{2})(t-s)^{3/2}+(Q+2\delta^{-1}\left|b(0)\right|^{2})(t-s).$
Taking limits as $m\rightarrow\infty$ we obtain
$\left\|\theta_{s,t}(\mu)\right\|_{2}^{2}\leq(1-2(\overline{b}-Q-\delta))(t-s))\left\|\mu\right\|_{2}^{2}+C(1+\left\|\mu\right\|_{2}^{2})(t-s)^{3/2}+(Q+2\delta^{-1}\left|b(0)\right|^{2})(t-s).$
We remark that the difference in the value of $\varepsilon^{\prime}$ between (B.5) and the result in Lemma B.1 is due to the fact that the above proof uses the modulus of continuity of the underlying process. This difference will also appear in the rest of the proof.
Step 3: Proof of the coupling property (2.22).
Lemma B.6.
A We assume that assumptions (5.35) and (5.36) hold with $Q<\overline{b}.$ Then, for every $b_{*}<2(\overline{b}-Q)$ there exists a constant $C$ such that, if $0\leq t-s\leq 1,$ then
${\mathbb{E}}\left|x_{s,t}^{1}(X^{1})-x_{s,t}^{2}(X^{2})\right|^{2}\leq{\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}(1-b_{*}(t-s))+C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(t-s)^{3/2}.$ (B.20)
In particular, $\theta_{s,t}$ and $\Theta_{s,t}$ are $(2,b_{*},1/2)$-coupled.
B Moreover $\theta$ is $(2,b_{*},1)$-self-coupled for some $h_{0}>0$ and more precisely
$W^{2}_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))\leq(1-b_{*}(t-s))W^{2}_{2}(\mu^{1},\mu^{2}){+C(1+\|\mu_{1}\|^{2}_{2}+\|\mu_{2}\|^{2})(t-s)^{2}}.$ (B.21)
Remark B.7.
When dealing with $W_{2}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))$ no error terms of the type ${\mathbb{E}}\left|b(x_{s,r}^{2}(X^{2}))-b(X^{2})\right|^{2}+{\mathbb{E}}\left|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})\right|^{2}$ appear in the proof. This is the reason why the term $C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(t-s)^{3/2}$ does not appear in the final result in comparison with (B.20).
Proof.
As in the previous proof we simplify the notation using $(x_{s,t}^{1},x_{s,t}^{2})=(x_{s,t}^{1}(X^{1}),x_{s,t}^{2}(X^{2}))$. Proof of A. The process $y_{s,t}=x_{s,t}^{1}-x_{s,t}^{2}$ satisfies
$y_{s,t}=X^{1}-X^{2}+\int_{s}^{r}\Delta b(r)dr+\int_{s}^{t}\int_{E}\int_{0}^{1}\Delta c(r,w,z)d\widetilde{N}(w,z,r),$
with
$\Delta b(r)$ $=$ $b(x_{s,\eta(r)}^{1})-b(X^{2})$
$\Delta c(r,w,z)$ $=$ $c(\tau_{s,\eta(r)}^{1}(w),z,x_{s,\eta(r)}^{1})-c(\tau_{s,\eta(r)}^{2}(w),z,X^{2})$
From (5.35) and using Lemma B.5, we get for every $\delta>0$
${\mathbb{E}}\left\langle y_{s,r},\Delta b(r)\right\rangle$ $\leq$ $-\overline{b}{\mathbb{E}}(|y_{s,r}|^{2})+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})\right\rangle+{\mathbb{E}}\left\langle y_{s,r},b(x_{s,r}^{2})-b(X^{2})\right\rangle$
$\leq$ $-\overline{b}{\mathbb{E}}(|y_{s,r}|^{2})+2\delta{\mathbb{E}}(|y_{s,r}|^{2})$
$+{\delta^{-1}}\left({\mathbb{E}}\left(|b(x_{s,r}^{2})-b(X^{2})|^{2}\right)+{\mathbb{E}}\left(|b(x_{s,\eta(r)}^{1})-b(x_{s,r}^{1})|^{2}\right)\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}\left|y_{s,r}\right|^{2}+{\delta^{-1}}C_{b}\left({\mathbb{E}}\left|x_{s,r}^{2}-X^{2}\right|^{2}+{\mathbb{E}}\left|x_{s,\eta(r)}^{1}-x_{s,r}^{1}\right|^{2}\right)$
$\leq$ $(2\delta-\overline{b}){\mathbb{E}}\left|y_{s,r}\right|^{2}+{\delta^{-1}}C_{b}C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s).$
Using the inequality $\left|\left|a\right|^{2}-\left|b\right|^{2}\right|\leq\left|a-b\right|(\left|a\right|+\left|b\right|)$, Cauchy-Schwarz inequality and (B.17), we have
$\left|{\mathbb{E}}\left|y_{s,r}\right|^{2}-{\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}\right|\leq C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s)^{1/2}.$ (B.22)
Combining the two last inequalities, we finally get for $r\leq s+1$ and a constant $C>0$ which depends on $\delta$
${\mathbb{E}}\left\langle y_{s,r},\Delta b(r)\right\rangle\leq(2\delta-\overline{b}){\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}+C(1+{\mathbb{E}}\left|X^{1}\right|^{2}+{\mathbb{E}}\left|X^{2}\right|^{2})(r-s)^{\frac{1}{2}}$
We now focus on the jump term and have by (5.36)
$\left|\Delta c(r,w,z)\right|^{2}\leq\overline{c}^{2}(z)(|\tau_{s,\eta(r)}^{1}(w)-\tau_{s,\eta(r)}^{2}(w)|^{2}+|x_{s,\eta(r)}^{1}(X^{1})-X^{2}|^{2}).$
By (B.17), ${\mathbb{E}}\left|x_{s,r}^{2}-X^{2}\right|^{2}\leq C(1+{\mathbb{E}}\left|X^{2}\right|)(r-s).$ We also recall that by the very definition of $\tau$, the inequality $2xy\leq\delta x^{2}+\delta^{-1}y^{2}$ and (B.17), we have
$\int_{0}^{1}\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}dw$ $=$ $W_{2}^{2}(\Theta^{m}_{s,r}(\mu^{1}),\mu^{2})\leq{\mathbb{E}}\left(|x_{s,r}^{1}(X^{1})-X^{2}|^{2}\right)$
$\leq$ $(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+(1+{\delta}^{-1}){\mathbb{E}}\left(|x_{s,r}^{1}-X^{1}|^{2}\right)$
$\leq$ $(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+C(1+{\delta}^{-1})(1+{\mathbb{E}}(|X^{1}|^{2}))(t-s).$
We conclude that
${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\nu(dz)dr$ (B.23)
$\leq$ ${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\overline{c}^{2}(z)(\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}+\left|x_{s,r}^{1}-X^{2}\right|^{2})dw\nu(dz)dr$
$\leq$ $2Q(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+C(1+{\delta}^{-1})(1+{\mathbb{E}}(|X^{1}|^{2}))(t-s).$
Then, the above estimates in (B.22) and (B.23) allow us to take expectations in the Itô’s formula for the function $x\rightarrow\left|x\right|^{2}$. We also use the identity $\left|y+\Delta c\right|^{2}-\left|y\right|^{2}-2\left\langle y,\Delta c\right\rangle=\left|\Delta c\right|^{2},$ so that we obtain
${\mathbb{E}}(|y_{s,t}|^{2})$ $=$ ${\mathbb{E}}(|X^{1}-X^{2}|^{2})+2{\mathbb{E}}\int_{s}^{t}\left\langle y_{s,r},\Delta b(r)\right\rangle dr+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\nu(dz)dr.$
$\leq$ ${\mathbb{E}}(|X^{1}-X^{2}|^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))+C(1+{\mathbb{E}}(|X^{1}|^{2})+{\mathbb{E}}(|X^{2}|^{2}))(t-s)^{3/2}$
so (B.20) is proved by choosing $\delta$ small enough. In particular, since ${\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}=W_{2}^{2}(\mu^{1},\mu^{2}),$ this gives
$W_{2}^{2}(\Theta_{s,t}^{m}(\mu^{1}),\Theta_{s,t}(\mu^{2}))$
$\leq$ ${\mathbb{E}}\left(|x_{s,t}^{1}-x_{s,t}^{2}|^{2}\right)={\mathbb{E}}(|y_{s,t}|^{2})$
$\leq$ $W_{2}^{2}(\mu^{1},\mu^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))+C(1+\left\|\mu^{1}\right\|_{2}^{2}+\left\|\mu^{2}\right\|_{2}^{2})(t-s)^{3/2}$
Now, using (5.39) and taking the limit as $m\rightarrow\infty$ we get
$W_{2}^{2}(\theta_{s,t}(\mu^{1}),\Theta_{s,t}(\mu^{2}))\leq W_{2}^{2}(\mu^{1},\mu^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))$
$+C(1+\left\|\mu^{1}\right\|_{2}^{2}+\left\|\mu^{2}\right\|_{2}^{2})(t-s)^{3/2}.$
We get that $\theta$ and $\Theta$ are $(b_{*},1,1/2)$-coupled for any $0<b_{*}<2(\bar{b}-Q)$.
The proof of B is analogue but simpler. ∎